### v1

Abstract
This paper presents a method for accurately predicting the full distribution of stock returns, given a comprehensive set of 194 stock characteristics and market variables. Such distributions, learned from rich data using a machine learning algorithm, are not constrained by restrictive model assumptions and allow the exploration of non-Gaussian, heavy-tailed data and their non-linear interactions. The method uses a two-stage quantile neural network combined with spline interpolation. The results show that the proposed approach outperforms alternative models in terms of out-of-sample losses. Furthermore, we show that the moments derived from such distributions can be useful as alternative empirical estimates in many cases, including mean estimation and forecasting. Finally, we examine the relationship between cross-sectional returns and several distributional characteristics. The results are robust to a wide range of US and international data.
Keywords: Distribution forecasting, Cross-section of stock returns, Anomalies, Quantile forecasting, Machine learning, Neural networks, Splines, International markets
JEL classification: C45, C53, C55, G12, G15, G17
1 Introduction
Despite considerable scepticism that returns are inherently difficult to predict (Stock and Watson, 2017), a substantial body of evidence has emerged over several decades showing that the first two moments of stock return distributions can be predicted to some extent using economic variables (Ang et al., 2006; Ang and Bekaert, 2007; McLean and Pontiff, 2016). As a practical matter, mean-variance analysis provides investors with a framework for making informed choices if they accept the constraints imposed by assumptions such as multivariate normality of stock returns or a quadratic utility function. At the same time, economic agents engaged in a wide range of activities, including risk management, portfolio selection, derivatives pricing and asset pricing, have a much more complex set of preferences. They are endowed with asymmetric utilities such as disappointment aversion (Gul, 1991), general preferences (Kimball, 1993), quantile preferences (Manski, 1988) or prospect theory (Kahneman and Tversky, 1979), which require distributional information beyond the first two moments. A complete distribution forecast is necessary to fully characterise the uncertain future evolution of returns for such decision-makers. A major challenge is to develop effective forecasting techniques for a return distribution, especially given the limitations of available data.
Forecasting stock return distributions is challenging due to a number of factors, including low signal-to-noise ratio, the presence of fat tails, volatility clustering and the potentially high-dimensional nature of the relevant information sets. Worryingly, much of the existing work relies on restrictive models and assumptions, does not take into account the possible non-linear interaction of a large number of variables, and does not take into account well-known features of the data, such as non-Gaussianity and heavy tails. With rapid improvements in the accessibility and availability of large datasets and computer science algorithms, it is tempting to ask whether the problem can be significantly improved using methods that focus on learning patterns from data. In line with recent efforts to move away from exclusive reliance on models to machine learning approaches (Athey and Imbens, 2019; Mullainathan and Spiess, 2017), we propose to use machine learning to predict the full distribution of out-of-sample stock returns. We provide ample evidence that such an approach makes sense using a wide range of US and international stocks.
Our contributions can be divided into two distinct areas. First, we present a novel distribution forecasting method based on the combination of a two-stage quantile neural network with cubic B-splines interpolation. The two-stage quantile neural network uses a substantial number of stock characteristics and market variables in two sub-networks to predict a set of $\tau$-quantiles of returns, which are then interpolated to obtain the full distribution. This crucial step allows the derivation and use of key moments of the distribution, including mean, volatility, skewness and kurtosis. The model has been trained on US data to demonstrate out-of-sample performance from both a time perspective, using a rolling window in the US, and a cross-sectional perspective on international stocks. The performance of the proposed method is compared with that of traditional parametric models and other neural network models. The results show that the proposed approach outperforms the benchmark models in terms of both out-of-sample average quantile loss and higher out-of-sample $R^{2}$ for mean forecasts, as well as lower out-of-sample error in variance forecasts. It is crucial to highlight that utilising information from the entire distribution enhances the forecasting accuracy of the mean and variance.
Second, the paper examines how different distributional characteristics are reflected in the cross-section of stock returns. A comprehensive dataset covering both US and international stocks is used in this study, with results presented for individual regions and for the full and liquid samples. In particular, we examine the influence of individual $\tau$-quantiles as well as mean, volatility, skewness and kurtosis derived from the predicted distribution. The profitability of long-short portfolios constructed on the basis of $\tau$-quantile forecasts of individual stocks is analysed. It is shown that central $\tau$-quantiles are priced into the cross-section of stock returns. There is a significant decrease in the profitability of long-short portfolios based on $\tau$-quantiles as we move from the centre of the distribution towards the tails. We show that the mean obtained by numerical integration of the interpolated distribution forecast leads to significantly higher out-of-sample long-short portfolio returns than the mean directly predicted by the competing models. In addition, we show that no other moment studied is significantly monotonically priced in the cross-section of stock returns. This provides additional support for previously conflicting results on the role of variance, skewness and kurtosis in asset pricing.
In order to illustrate the efficacy of our methodology, we present the forecasted probability function obtained using our methods on the case of the Apple stock. Figure 1 illustrates the evolution of the forecasted distribution of the 22-day-ahead returns over the period 1996-2020. The distribution is obtained without any model assumptions, utilising a substantial number of characteristics. Yet, it captures the dynamics precisely.
Figure 1: Example of forecasted probability density function over time. This figure displays the forecasted probability density function for out-of-sample 22 days-ahead returns of an Apple stock (CUSIP 03783310) generated using our two-stage quantile neural network model.
In our methodology, we prioritise the forecasting of quantiles and subsequently approximate the distribution function rather than directly estimating a parametric distribution. In light of the stylised facts of stock returns, including the presence of fat tails, asymmetry and a domain extending from $-1$ to +$\infty$+ ∞, it is challenging to specify a distributional form that accurately captures these characteristics. For example, the normal distribution is unable to capture the fat tails, the Student-t distribution is unable to capture the asymmetry, and the skewed Student-t distribution is unable to capture the domain restriction. Conversely, our approach approximates the distribution function in a data-driven manner, eschewing the imposition of strong parametric assumptions. Furthermore, our method is capable of estimating discrete probabilities, such as the probability mass at -1, which is of particular importance for long-horizon return forecasts. The selected methodology yields intriguing outcomes, such as the predicted discrete probability mass at 00 for certain stocks, which would be unattainable through the parametric distribution function.
The empirical results presented in this study are based on a comprehensive analysis of both U.S. and international stock data spanning the period from 1995 to 2018. To ensure the robustness of our results, we focus on the full sample as well as a restricted liquid sample. To demonstrate the generalisability of our distribution forecasting method, we train the model using only U.S. data and then apply it to international stocks and U.S. stocks in an out-of-sample manner. All requisite stock-level characteristics and market variable definitions are consistent across the United States, thereby eliminating the potential for country-specific variables (e.g., macroeconomic variables) or variables available only for a subset of stocks (e.g., options-based variables) to influence the model. This consistency allows for the transfer of predictability to the international domain.
By focusing on the aspect of predictability, we demonstrate that our distribution forecasting method exhibits a notable degree of outperformance in terms of average out-of-sample quantile loss when compared to forecasts generated by the traditional time series models as well as complex neural networks. Furthermore, the distribution forecast provides substantial benefits when forecasting the mean of stock returns. The mean forecast obtained by means of numerical integration of the interpolated distribution forecast gives rise to significantly higher out-of-sample $R^{2}$ values than the mean forecast derived directly from neural network models using mean squared error loss. With regard to volatility forecasting, the volatility forecast derived from numerical integration of the interpolated distribution forecast results in a reduction in the out-of-sample mean absolute deviation and root mean squared error in comparison to those forecasted from the GARCH model. These findings are consistent across all regions and both full and liquid samples.
In light of the distribution forecasting capability of our model, we undertake a further empirical investigation into the role of different distributional characteristics in an asset pricing application. This study examines the profitability of long-short portfolios formed based on $\tau$-quantile forecasts of individual stocks, as well as the roles of mean, volatility, skewness, and kurtosis. Our findings indicate that, among all the quantiles, only the central quantiles are reflected in the cross-section of stock returns. Furthermore, we observe a notable decline in the profitability of long-short portfolios formed on these quantiles as we move away from the central region of the distribution towards the tails. Furthermore, we examine the profitability of long-short portfolios based on the mean versus the median forecast. Our findings indicate that the mean forecast leads to significantly higher out-of-sample long-short portfolio returns on the full sample. However, this is not the case on the liquid sample, which highlights the importance of underlying distributional asymmetry. Finally, we demonstrate that no other studied moment is priced in the cross-section of stock returns, providing further evidence to resolve the previously observed conflicting results on the roles of variance, skewness, and kurtosis in asset pricing. The following section will present the results of the analysis.
1.1 Relation of our work to the literature
Our work relates to several strands of the literature. We contribute to stock return forecasting, which is a central topic in the asset pricing literature with a long history. In recent years, machine learning methods have become a tool of choice in this literature due to their ability to capture complex non-linear relationships, handle high-dimensional data with low signal-to-noise ratios, and provide superior sample-to-sample performance compared to traditional approaches and superior out-of-sample performance compared to more traditional approaches (Gu et al., 2020). For example, Bryzgalova et al. (2020) shows how to build better test assets by creating managed portfolios using decision trees, Dong et al. (2022) links the cross-sectional predictability of stock returns with the time-series predictability of aggregate market returns by using anomaly portfolios to predict market returns, Gu et al. (2021) uses autoencoders to build a factor asset pricing model that allows for non-linear factor loadings. Successful applications of machine learning techniques have also been demonstrated in other domains and asset classes, such as the prediction of bond returns by Bianchi et al. (2021), the ex-ante identification of good and bad performing mutual funds by Kaniel et al. (2023), the prediction of option returns by Bali et al. (2023), or the use of convolutional neural networks to use price images to predict stock returns by Jiang et al. (2023). Furthermore, the robustness of many machine learning applications in stock return forecasting has been confirmed in an international context by Azevedo et al. (2023), Cakici et al. (2023) and Tobek and Hronec (2021).
The literature on forecasting the distribution of stock returns is more limited. Our work is most closely related to Liu (2023) in terms of forecasting the quantiles of future stock returns in the cross-section. Liu (2023) follows Gu et al. (2020) and compares a variety of linear models and machine learning methods in a quantile regression setting. [^1] Models compared include linear quantile regression, elastic net quantile regression, quantile random forests, quantile gradient boosted regression trees, and feed-forward neural networks. In terms of out-of-sample prediction performance, he shows that quantile neural networks significantly outperform other models. He introduces a quantile-based risk premium measure as a weighted average of five predicted quantiles, where the weights correspond to an area under an empirical step function. He also constructs a robust measure of volatility (interquantile range) and skewness from the $25\%$, $50\%$ and $70\%$ quantile forecasts, documenting that skewness is priced into the cross-section of stock returns. Our paper differs from Liu (2023) in several ways.
We present a novel neural network architecture for quantile forecasting and show how a list of quantile forecasts can be interpolated into a dense representation of the distribution function. Given our denser representation of the distribution, we derive moments using numerical integration, which provides a more accurate representation of the moments than directly using robust quantile-based measures. We also show how to adjust the predicted higher moments to account for the lack of extreme observations in the tails of the distribution. In addition, we provide international evidence for models trained on US data only and show that our approach generalises well to international stocks. Our results are consistent with the importance of obtaining the mean prediction from the distribution [^2] Even if the implementation differs. rather than directly using the mean squared error; however, we find no evidence that skewness is priced into the cross-section of stock returns, either in the US or internationally. We also examine the role of individual $\tau$-quantiles in the cross-section of stock returns.
Another related paper is Yang et al. (2024), which builds on Gu et al. (2021) by using a conditional autoencoder model to approximate the distribution of future stock returns via a step function using a conditional quantile variational autoencoder (CQVAE). The authors show that the CQVAE model effectively captures the non-linear dependencies between stock returns and latent factors, leading to superior out-of-sample forecasting performance compared to traditional linear models. They also show that the mean returns estimated from the step-function approximation of the distribution function are more accurate and lead to better portfolio performance. While both our paper and Yang et al. (2024) use machine learning to predict the distribution of stock returns, there are significant differences. The CQVAE model uses latent factors learned by a variational autoencoder, while our approach uses stock characteristics and market variables directly as predictors. We go further by deriving key distributional moments such as mean, volatility, skewness and kurtosis through numerical integration, in contrast to the CQVAE model’s focus on using quantile forecasts primarily for mean estimation. In addition, we validate the generalisability and robustness of our model using both US and international stock data, demonstrating its broad applicability, whereas Yang et al. (2024) focuses solely on US stocks.
The distributions of individual stock returns are relevant to the asset pricing literature, as can be seen from the from the number of anomalies based on different distribution characteristics, see Hou et al. (2020). However, these anomalies are typically based on historical sample estimates rather than out-of-sample forecasts, leading to significant estimation errors, especially when the persistence of tail-based distributional of tail-based distributional properties is low. For example, estimating the third and fourth moments rather than the first two moments Kim and White (2004). A promising alternative to distributional forecasting is the use of options to derive implied moments (An et al. (2014), Huang and Li (2019), Alexiou and Rompolis (2022)), although this method is limited by the coverage and liquidity of the and liquidity of options markets, especially outside the US. Instead of relying on historical sample estimates or option data, we use distributional characteristics derived from our out-of-sample distribution forecasts.
A large body of literature examines how distributional characteristics affect the cross-sectional distribution of stock returns. of stock returns. There is a consensus on the ambiguous relationship between realised volatility and future stock returns. and future stock returns (Bollerslev et al. (2020), Ang et al. (2006)). We contribute further evidence on the lack of a clear relationship between forecast volatility and future stock returns. stock returns. A similar lack of documented relationships also exists for kurtosis, although the the literature here is less extensive (Bollerslev et al., 2020).
The role of skewness, on the other hand, presents a more complex picture. Skewness is crucial in understanding individual stock returns, as Bessembinder (2018) notes that the positive skewness of individual stock returns fundamentally explains why a small fraction of stocks generate long-term stock market wealth. A series of papers, extensively reviewed by Bali et al. (2016), examine the risk premium associated with the cross-sectional skewness of stock returns. of stock returns.
In addition to asset pricing, the importance of predicting distributions is evident in risk management, where the left tail of the distribution is of particular interest. Value at Risk (VaR) is a standard risk management tool and corresponds to the $\tau$ quantile of the conditional distribution of an asset’s return. Various approaches to VaR estimation, ranging from non-parametric (historical simulation) to semi-parametric to parametric, have a long history in the literature. Unlike non-parametric estimation, parametric VaR estimation requires assumptions about specific return distributions and model dynamics. The foundation of this approach is based on the GARCH volatility model of Engle (1982) and Bollerslev (1986), which is typically used together with the assumption that the residuals are distributed according to a specific distribution, typically normal or Student-t. Among semi-parametric estimation approaches, a prominent example is the CaViaR model of Engle and Manganelli (2004), which uses quantile regression to directly estimate the model dynamics without making further assumptions about the error distribution. See Nieto and Ruiz (2016) for a comprehensive review of VaR estimation methods.
While our work and the VaR estimation literature are clearly related in that both rely on accurate quantile forecasts, there are also obvious differences. First, we focus on the entire cross-section of stock returns, while the VaR estimation literature tends to work with a single asset, usually representing the entire portfolio. This means that we can rely on a rich set of information in the form of stock-level characteristics, whereas the VaR estimation literature must rely mainly on the historical returns of the asset itself. Second, our ultimate goal is to forecast the full distribution of stock returns, not just specific $5\%$ or $1\%$ quantiles. As a result, we can derive any distributional risk measure, such as conditional value at risk (CVaR), skewness or kurtosis.
The rest of the paper is organised as follows. Section 2 introduces our distribution forecasting method, including quantile neural networks and subsequent interpolation of the quantiles into a full distribution function using splines. Section 3 presents the empirical results, consisting of an out-of-sample assessment of predictability and an analysis of the asset pricing implications of various distributional features.
2 Two-stage quantile neural networks
Let the returns $r_{i,t}$ collected over $t=1,\dots,T$ months and $i=1,\dots,N_{t}$ individual stocks be random variables with a distribution function [^3] In the case of the forecast distribution of stock returns, $Q_{r_{i,t}}(0)=-1$ and $Q_{r_{i,t}}(1)=\infty$. $F(r_{i,t})$. We are interested in forecasting a set of multiple $\tau$ quantiles $Q_{r_{i,t+1}}^{\mathcal{T}}=\Big{\{}Q_{r_{i,t+1}}(\tau_{1}|\mathcal{F}_{t}),\ldots,Q_{r_{i,t+1}}(\tau_{K}|\mathcal{F}_{t})\Big{\}}$ for $\forall\tau\in\mathcal{T}=\{\tau_{1},\ldots,\tau_{K}\}$ where $Q_{r_{i,t}}(\tau\big{|}\mathcal{F}_{t})=\inf\{r_{i,t}\in\mathbb{R}:F(r_{i,t})\geq\tau|\mathcal{F}_{t}\}$ of a stock return at time $t+1$ is conditional on the information set $\mathcal{F}_{t}$ available at time $t$.
Forecasting multiple $\tau$-quantiles is possible with a multi-output feed-forward neural network model based on aggregating individual $\tau$-quantile losses into a single loss function (as described later in this section). In addition to being computationally more efficient than training individual models for each $\tau$, joint training for all $\tau$ quantiles allows the model to capture the relationship across the entire distribution. As a side effect, even without introducing penalisation for quantile crossing [^4] Quantile crossing refers to the situation where $Q(\tau_{l})>Q(\tau_{h})$ when $\tau_{l}<\tau_{h}$, violating the quantile monotonicity condition., the model naturally captures the monotonicity requirement of quantile predictions since there are no quantiles that are crossed in our quantile forecasts.
2.1 Network architecture
While neural networks have been successfully used in stock return forecasting (Gu et al., 2020), quantile forecasting is challenging because we need to capture both cross-sectional variation across stocks and market-wide variation over time. One way to address this challenge is to use cross-sectional standardisation of returns to remove the time series noise. This is possible from an asset pricing perspective, where the goal is to rank individual stocks cross-sectionally to create long-short portfolios. However, since our goal is to forecast quantiles of future raw (non-standardised) stock returns, we propose a two-stage neural network architecture that exploits the predictability of standardised returns in the first stage and generates quantile forecasts of true (non-standardised) stock returns in the second stage. In this way, we allow the neural networks to exploit the information in standardised returns and to adjust for expected market volatility.
More specifically, a set of quantiles of stock returns are modelled using a large set of variables, including stock-level characteristics (e.g., size, value, or momentum) stored in a vector $x_{i,t}$, as well as market variables (e.g., market volatility) stored in a vector $z_{t}$, both of which proxy the information available at period $t$. We use both raw and standardised returns to address the discussion above, and we standardise the return of a stock by dividing the raw stock returns by the cross-sectional average of the stock-level volatilities at period $t$, such that
$\widetilde{r}_{i,t+1}=\frac{r_{i,t+1}}{\overline{\sigma}_{t}},$ (1)
where $\overline{\sigma}_{t}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\sigma_{i,{t}}$ is a cross-sectional average of stock-level volatilities, $\sigma_{i,t}$ is a stock-level volatility estimate of stock $i$ at period $t$, and $N_{t}$ is the number of stocks in the panel at period $t$. Ideally, we would use $\overline{\sigma}_{t+1}$ instead of $\overline{\sigma}_{t+1}$ to rescale the standardised quantile forecasts back to the raw quantile forecasts of stock returns, but this is not known at period $t$ when the forecast is made. Therefore, we use the cross-sectional average of stock-level volatilities at period $t$ instead of the future value of $\overline{\sigma}_{t+1}$ to avoid look-ahead bias. $\sigma_{i,t}$ is estimated using an exponentially moving weighted average of the squared returns of stock $i$ with a decay factor of $0.94$ and is one of the input variables to the model. [^5] The 0.94 parameter is widely used in both academia and the financial industry, an example being the RiskMetrics GARCH specification. See Appendix C.1 for a visualisation of $\overline{\sigma}$ over time in the US.
The idea behind standardising individual stock returns using the cross-sectional average of stock-level volatilities rather than the stock volatility itself is to reduce the noise from individual stock volatility estimates. For the same noise-reducing reason, we standardise stock returns only by dividing by the average volatility and and not by subtracting the average return, either stock or market, because average returns are not persistent.
A stock return quantile is then modelled by the two-stage quantile neural network with two sub-networks as
$Q_{r_{i,t+1}}^{\mathcal{T}}$ $=$ $\Big{\{}Q_{r_{i,t+1}}(\tau_{1}|x_{t},z_{t},\overline{\sigma}_{t}),\ldots,Q_{r_{i,t+1}}(\tau_{K}|x_{t},z_{t},\overline{\sigma}_{t})\Big{\}}$ (2)
$=$ $\Bigg{[}\Big{(}\underbrace{g^{(L)}_{W_{S_{1}}^{(L)},b_{S_{1}}^{(L)}}\circ\ldots\circ g^{(1)}_{W_{S_{1}}^{(1)},b_{S_{1}}^{(1)}}\left(x_{t}\right)}_{\text{standardized $\tau$ quantiles sub-network}}\Big{)}\times\overline{\sigma}_{t}\Bigg{]}\times\Big{(}\underbrace{g^{(2)}_{W_{S_{2}}^{(2)},b_{S_{2}}^{(2)}}\circ g^{(1)}_{W_{S_{2}}^{(1)},b_{S_{2}}^{(1)}}\left(z_{t}\right)}_{\text{market volatility sub-network}}\Big{)}$
$=$ [$\underbrace{\Bigg{[}\Big{\{}\underbrace{Q_{\widetilde{r}_{i,t+1}}(\tau_{1}|x_{t}),\ldots,Q_{\widetilde{r}_{i,t+1}}(\tau_{K}|x_{t})}_{\text{Stage I}}\Big{\}}\times\overline{\sigma}_{t}\Bigg{]}\times\widehat{\sigma}_{t}^{M}}_{\text{Stage II}},$
where $W_{S}=\left(W_{S}^{(1)},\ldots,W_{S}^{(L)}\right)$ and $b_{S}=\left(b_{S}^{(1)},\ldots,b_{S}^{(L)}\right)$ are weight matrices and bias vectors from a stage $S$. Each weight matrix $W_{S}^{(\ell)}\in\mathbb{R}^{m\times n}$ contains $m$ neurons as $n$ column vectors $W_{S}^{(\ell)}=[w_{\cdot,1}^{(\ell)},\ldots,w_{\cdot,n}^{(\ell)}]$, and $b_{S}^{(\ell)}$ are thresholds or activation levels that contribute to the output of a hidden layer that allows the function to be shifted.
It is important to note that, in contrast to the literature, we consider a multi-output (deep) neural network to characterise the collection of quantiles. Our network has two subnetworks, namely a standardised $\tau$ quantile network and a market volatility network in which $l\in{1,...,L}$ hidden layers transform the input data into a chain using a collection of nonlinear activation functions $g_{S}^{(1)},\ldots,g_{S}^{(L)}$. An activation function, $g^{(\ell)}_{W_{S}^{(\ell)},b_{S}^{(\ell)}}$, is used as a
$g^{(\ell)}_{W_{S}^{(\ell)},b_{S}^{(\ell)}}(u):=g_{\ell}\left(W_{S}^{(\ell)}u+b_{S}^{(\ell)}\right)=g_{\ell}\left(\sum_{i=1}^{m}W^{(\ell)}_{i,S}u+b_{i,S}^{(\ell)}\right)$
is rectified linear units $g_{\ell}(u)=\max\{u,0\}$, or $g_{\ell,S}(u)=\tanh(u)$.
$\overline{\sigma}_{t}$ $Q^{\mathcal{T}}_{\widetilde{r}_{i,t}}$ (37$\times$1) $\times$$\times$$\times$(128$\times$2)(128$\times$2) Bottleneck (4$\times$1) $x_{t}$ (176$\times$1) $\times$$\times$$\times$ $z_{t}$ (18$\times$1) (8$\times$1) $\widehat{\sigma}_{t}^{M}$ (1$\times$1) $Q^{\mathcal{T}}_{r_{i,t}}$ (37$\times$1) Stage IStage IIStandardized $\tau$-quantiles sub-networkMarket volatility sub-network
Figure 2: Two-stage neural network architecture. Stage I generates cross-sectionally standardised $\tau$-quantile forecasts, $Q^{\mathcal{T}}_{\widetilde{r}}$, as output. Cross-sectional standardisation is performed by dividing individual future stock returns $r_{i,t+1}$ (labels) by the cross-sectional average of the individual stock standard deviation estimates, $\overline{\sigma}_{t+1}$. The standard deviation estimate is an exponential weighted moving average of volatility with a smoothing factor of $0.94$. Step II uses $\overline{\sigma}_{t}$, i.e. based on data from the previous period, to rescale the standardised $\tau$-quantile forecasts generated in Stage I back to the original scale. The rescaling takes place at the multiplication nodes marked with $\times$, representing the multiplication operation between the between the inputs of the node. The market volatility part then generates forecasts of the next period’s market volatility in order to to further refine the $\tau$-quantile forecasts produced as the output of stage II. The forward pass generates a tensor of form $(B,37,2)$, where $B$ is the number of observations in the batch, 37 is the number of predicted quantiles, and 2 corresponds to the standardised and raw versions.
Figure 2 details the architecture of the two-stage quantile neural network model defined above. This is our main model used for quantile forecasting, and we refer to it as the two-stage model throughout the paper. The first stage of the model is used to forecast quantiles of cross-sectionally standardised next month stock returns $Q^{\mathcal{T}}_{\widetilde{r}_{i,t}}$. We use $37$ $\tau$s in the set $\mathcal{T}$ ranging from the extreme left tail of the distribution (0.00005) to the extreme right tail of the distribution (0.99995). The second stage of the model generates quantile forecasts of the next month’s raw stock returns $Q_{r_{i,t}}(\tau)$ for $\tau\in\mathcal{T}$. This stage consists of three steps. First, the standardised quantile forecasts $Q^{\mathcal{T}}_{\widetilde{r}_{i,t}}$ from the first stage and the cross-sectional average of stock-level volatilities $\overline{\sigma}_{t}$ are used as inputs to rescale the standardised quantile forecasts back to the raw stock return scale $Q_{r_{i,t}}(\tau)=Q_{\widetilde{r}_{i,t}}(\tau)\cdot\overline{\sigma}_{t}$. Second, the market volatility subnetwork generates a second scaling factor to reintroduce the time-series market-wide volatility component. This market volatility scaling factor is the same for all stocks in a given period $t$ and is used to refine the quantile forecasts rescaled in the first step. Finally, both the standardised and raw quantile forecasts are clipped at -100% during the forward pass of the neural network.
In addition to the use of two stages, another key difference from standard feedforward neural network architecture is the inclusion of a bottleneck layer. Inspired by the autoencoder architecture, the bottleneck layer aims to capture the distribution hyperparameters. This is achieved by forcing the network to learn a compressed representation, as the bottleneck layer is limited to only four nodes.
The two-stage neural network takes three sets of inputs. The first set consists of 176 features representing stock-level characteristics based on 153 anomalies used in Tobek and Hronec (2021), 18 stock-level volatility estimates, and five cross-sectional averages of stock-level mean estimates. The second set of inputs consists of cross-sectional averages of the stock-level volatility characteristics used in the first set, i.e. 18 stock-level volatility estimates. The third input is a cross-sectional average of stock-level volatilities $\overline{\sigma}_{t}$ used for rescaling purposes. A detailed description of all features can be found in Section 3.1.1.
2.2 Training with quantile loss function
In order to jointly predict multiple $\tau$-quantiles using a two-stage neural network, $\tau$-quantile losses for individual $\tau$s must be aggregated into a single loss. The two-stage quantile neural network model also requires both standardised and raw stock returns as labels, so the aggregated quantile loss must take this into account. The loss function used during training is defined as a linear combination of the aggregated quantile losses from the first and second stage outputs, as shown in Equation 3. Both coefficients in the linear combination are equal, implying equal weighting of both stages in the loss function. [^6] This specification worked out of the box, and the hyperparameter was therefore not calibrated. In practice, the first stage has a much larger impact on the loss function because it focuses on rescaled raw returns with a scale close to one, while the second stage focuses on raw returns with a scale of about 0.1. The greater implicit focus on the first stage is desirable, as this is where most of the learning takes place, with the second stage only fine-tuning the results of the first stage. The aggregated multi-$\tau$-quantile loss used in the estimation is defined as
$\mathcal{L}^{\mathcal{T}}=\frac{1}{B}\frac{1}{K}\sum_{\tau\in\mathcal{T}}\sum_{i=1}^{B}\left(\rho_{\tau}\left(r_{i,t}-\widehat{Q}_{r_{i,t}}(\tau)\right)+\rho_{\tau}\left(\widetilde{r}_{i,t}-\widehat{Q}_{\widetilde{r}_{i,t}}(\tau)\right)\right)$ (3)
where $\widehat{Q}_{r_{i,t}}(\tau)$ is the forecast $\tau$-quantile of the raw return of stock $i$ at period $t$, $\widehat{Q}_{\widetilde{r}_{i,t}}(\tau)$ is the predicted $\tau$-quantile of the standardised return of stock $i$ at period $t$, $B$ is the lot size and $K=37$ is the number of $\tau$ in the $\mathcal{T}$ set. An essential part of the loss is $\rho_{\tau}$, a quantile loss function defined as a piecewise linear function $\rho_{\tau}(\xi)=\mathds{1}_{(\xi\geq 0)}\tau\xi+\mathds{1}_{(\xi<0)}(\tau-1)\xi$ where $\xi=y-\widehat{Q}_{y}(\tau)$ is an error term between the true value and the predicted $\tau$ quantile, and $\mathds{1}_{(.)}$ is an indicator function. A single $\tau$ quantile loss calculated for multiple observations is then calculated as a simple arithmetic mean of the individual $\rho_{\tau}$ losses. The idea behind the quantile loss function is to penalise overestimation of the $\tau$ quantile by $\tau$ times the error if the error is positive and by $(\tau-1)$ times the error if the error is negative. For example, if $\tau=0.5$, the quantile loss function is the absolute value loss function, and overprediction and underprediction are penalised equally. If $\tau=0.05$, underprediction is penalised $19$-times $(0.95/0.05)$ more than overprediction, and minimising such a loss function leads to the estimation of the $0.05$ quantile.
This definition implies an equal weighting of individual $\tau$-quantile losses in the aggregate loss function. In our case, this means that large stocks have the same impact as small stocks and central quantiles have the same impact as tail quantiles.
From a numerical point of view, the quantile loss function is a piecewise linear function that is differentiable everywhere except at zero. Gradient-based optimisation algorithms require the loss function to be differentiable in order to update the model parameters. Automatic differentiation engines, such as PyTorch’s autograd (Paszke et al. (2019)) are able to deal with this problem by defining subgradients for non-differentiable points. A subgradient at a point is any slope of a line that lies below the function at that point and intersects the function at that point. For the quantile loss function, the subgradient at zero can be any value between the left derivative ($\tau-1$) and the right derivative ($\tau$), depending on the specific implementation of the quantile loss function.
To train the model, we minimise the loss function using stochastic gradient descent with mini-batches of size $B=8192$ observations. Specifically, we use the Adam optimiser (Kingma and Ba (2014)) with an initial learning rate of $0.0003$ and $(\beta_{1},\beta_{2})=(0.9,0.999)$. We adjust the number of epochs based on the number of observations available at each training period. The number of epochs is set to $100*(A/n)$, where $n$ is the number of observations available for training and $A$ is the adjustment constant equal to $3000000$ for the full sample and $1500000$ for the fluid sample, derived from the average number of observations in each sample over time. [^7] See Section 3.1 for the description of the data used. We use early stopping with a patience of 2 epochs, where 20% of the training data is used as validation to determine the stopping point. After early stopping, the model state with the lowest validation loss is taken as the final model. We use Leaky Relu (Maas et al. (2013)) as the activation function in all layers except the last layer in each subnetwork. We use batch normalisation (Ioffe and Szegedy (2015)) in all layers except the last layer in both the standardised $\tau$-quantiles subnetwork and the market volatility subnetwork. We apply dropout to all layers except the last layer in each subnetwork and use a rate of $0.2$ as a result of the hyperparameter search described in Appendix D. We also apply an L1 penalty to the weights of the first stage in each subnet. The L1 penalty for the first layer is $0.0001$, and for the second layer, it is $0.00001$. The first layer of the market volatility network is also regularised by an L2 penalty of $0.00001$. As an additional form of regularisation, we use an ensemble of 20 networks.
The model is trained from scratch for the first training period and then fine-tuned for subsequent training periods [^8] Initial weights are taken from the previous training period.. Fine-tuning greatly reduces the computational burden of having to re-estimate the model each period and is a natural choice given the large overlap of subsequent training samples with the widening estimation window approach. It is expected that one year of new data should not change the estimated parameters too much unless there is a novel high-information event such as the dot-com bubble or the global financial crisis, in which case fine-tuning requires more epochs to retrain.
2.3 Approximating probability density and its moments from quantiles using B-Splines
To approximate the probability density, let’s consider a grid of $\tau$ quantile forecasts $\{Q_{j}(\tau),\tau_{j}\}_{j=1}^{K}$, where $\tau_{j}$ is the probability and $Q_{j}(\tau)$ is the corresponding quantile value, containing $K$ grid points that can be used to obtain the cumulative distribution function $F:\mathbb{R}\rightarrow[0,1]$. Since stock returns range from $-1$ to $\infty$, from a practical perspective we are interested in $F:[-1,\infty)\rightarrow[0,1]$. To get a denser representation, we use cubic B-splines [^9] We use the splrep and BSpline functions from the Python scipy library, Virtanen et al. (2020). to interpolate between adjacent quantiles by constructing piecewise polynomial functions within each interval $[Q_{j},Q_{j+1}]$. Instead of interpolating the predicted quantiles directly, we create a new denser $\mathcal{T}^{d}$ grid with one hundred equally spaced points between every two original quantile prediction points, except the largest and smallest, i.e. the denser grid has $(K-3)\times 100$ values, and fit the cubic B-spline to approximate the cumulative distribution function. To derive the probability density function, the derivative of the spline for the base function is calculated at each interval to obtain the new spline function representing the probability density function.
Figure 3 shows an example of the interpolated cumulative distribution function on the left and the implied probability density function on the right. The figure is based on $\tau$-quantile forecasts of 22-day (monthly) returns for Microsoft in October 2008. The density function is relatively close to a Student’s t-distribution with noticeably large tails during the depths of the Great Financial Crisis.
Figure 3: Interpolated cumulative distribution function and probability density function. This figure shows the cumulative distribution function interpolated from the inverted quantile forecasts depicted as marks (left) and derived probability density function (right) for out-of-sample 22 days-ahead returns of Microsoft stock (CUSIP 59491810) in October 2008. Quantile forecasts used to obtain the cumulative distribution function are generated using the two-stage quantile neural network model.
As the interpolated density function does not cover the extreme tails of the distribution beyond the smallest and largest predicted quantile, these tails are modelled by discrete probabilities for values at the ends of the new $\mathcal{T}^{d}$ grid. We use discrete probabilities implied by the estimated spline function. The density function is truncated for returns less than -1, and discrete probability is then considered for $-1$ returns. Excluding the extreme tails from the density function results in a slight underestimation of the variance and a more pronounced underestimation of the kurtosis of the distribution. Inferring the distribution in the tails from empirical data is problematic, so we chose this solution to avoid the problem to some extent. In cases where B-spline interpolation does not produce plausible results, we use a linear B-spline function as a fallback (see Appendix E.1 for details). A complete algorithm for approximating the probability density function from the quantile forecasts is described in detail in Appendix E, Algorithm 1.
Importantly, the B-spline approximation allows us to obtain moments of the probability distribution function. We first derive the $z$th non-central moment of a density function $f(x)$
[…]
using numerical integration methods and the B-spline-based density function approximation obtained above. We use a simple tractable local linear approximation algorithm, which leads to trivial integrals of polynomial functions, and the complete algorithm is detailed in Appendix E, Algorithm 2. Further, using this approximation, we calculate central moments and rescale the moments by the integrated density function to ensure they are properly defined.
The estimated empirical quantiles from neural networks do not cover the full range of possible stock returns because it is not possible to estimate quantiles in the extreme tails of the distribution. The calculated central moments are, therefore, biased, with the bias being greater for higher moments as the tails of the distribution have a greater impact. It is, therefore, important to test how well the algorithm captures the true moments. To do this, we compare the true moments of a theoretical distribution with the output of the algorithm and propose an adjustment to reduce the measured bias. The appendix Section E.3 describes how to adjust the moments to account for bias.
3 Empirical results
3.1 Data and features
Our data cover U.S. and international equity markets for the periods 1963 to 2018 and 1990 to 2018, respectively. We use the merged CRSP/Compustat database from the Wharton Research Data Service (WRDS) for the US sample and LSEG Datastream101010formerly called Reuters or Refinitiv Datastream. for the international sample. Data are pre-processed using the methodology described in Tobek and Hronec (2021). There are four regions in the sample - US, Europe, Japan and Asia-Pacific, which also includes 23 developed countries. [^11] US, Austria, Belgium, Denmark, Finland, France, Germany, Greece, Ireland, Italy, Luxembourg, Netherlands, Norway, Portugal, Spain, Sweden, Switzerland, UK, Japan, Australia, New Zealand, Hong Kong and Singapore. We report our results for both the full sample and the liquid sample, which includes only the most liquid stocks. The liquid sample includes stocks that meet the following criteria: a price greater than $1$ (or $0.1$ for Asia Pacific) at the end of the previous month, a market capitalisation within the top 95%, and a trading volume within the top 95%, and a trading volume within the top 95% of the total dollar trading volume over the previous 12 months in each region. The full sample includes stocks with a price greater than $1$ (\$0.1 for Asia Pacific) at the end of the previous month. Market capitalisation and volume filters are not applied to the full sample. Micro-cap stocks in the full sample therefore account for only a small fraction of the capitalisation and capitalisation and trading volume of the total market. The average, minimum and maximum number of stocks in the cross-section in each region is shown in table Table 1.
Table 1: Number of stocks in the cross-section across the regions. This table shows the average, minimum and maximum number of stocks in the cross-section in each region for the full and liquid samples. The period covered is from 1973 to 2018 for the US sample and from 1990 to 2018 for the international sample.
Full Sample Liquid Sample
region USA Europe Japan Asia Pacific USA Europe Japan Asia Pacific
min 3552 5098 1919 1157 873 413 534 226
mean 5136 5746 3225 2834 1167 690 749 430
max 7384 6740 3795 4553 1734 1044 1079 712
3.1.1 Features
The two-stage quantile neural network model uses both stock characteristics and market variables aggregated from the stock-level variables as features. There are 176 stock characteristics consisting of 153 anomalies from Tobek and Hronec (2021) plus an additional 18 stock-level volatility estimates. The 23 market-level variables consist of 18 market volatility variables and 5 market mean variables. This brings the total number of features to 194.
The list of 153 anomalies is given in Appendix A and includes 93 fundamental, 49 market friction and 11 I/B/E/S anomalies, mostly covered by McLean and Pontiff (2016), Hou et al. (2020) and Harvey et al. (2016). Compared to the original data, we use weekly data during training. This means that The fundamental signals are updated every week with financial statement information from financial years ending at least 6 months earlier. Trade data information, such as market capitalisation, is taken as the most recent available. Missing characteristic values are imputed using cross-sectional medians.
Each sub-network in the two-stage quantile neural network model uses its own set of features. The standardized $\tau$-quantile sub-network uses 176 features consisting of 153 anomalies-based stock characteristics together with 18 stock-level volatilities and 5 market mean returns. The 18 stock-level volatilities are calculated using daily exponentially weighted moving averages (EWMA) with $\alpha\in\{0.8,0.9,0.94,0.96,0.98,0.99\}$ smoothing factors, EWMA of negative returns with $\alpha$ smoothing factors of $\{0.8,0.9,0.94\}$ to proxy asymmetric volatility, total volatilities calculated over the past 3, 6 and 12 months, and Parkinson (1980) high-low range-based volatility. All stock-level volatility estimates are scaled by their cross-sectional mean to remove the effects of changes in market-wide average volatility. This is done to make stock-level volatilities more consistent with the other stock-level variables, which are normalised to cross-sectional empirical quantiles. The 5 additional market variables used as inputs to the standardized $\tau$-quantile sub-network are computed as EWMAs of cross-sectional equal-weighted averages of individual stock returns $\mu_{t}=\alpha\cdot\left(1/N_{t}\sum_{i=1}^{N_{t}}r_{i,t-1}\right)+(1-\alpha)\cdot\mu_{t-1}$ over the number of stocks $N_{t}$ with smoothing parameters $\alpha\in\{0.9,0.94,0.96,0.99,0.999\}$. These variables measure the average historical market-wide return and should help in forecasting when there is some persistent time variance in expected market mean returns. We rescale the mean variables by the same scalar as the predicted returns to reflect the fact that their scale over time is relevant in the estimation.
The market volatility sub-network uses 18 cross-sectional averages of stock-level volatilities described above. These serve as a proxy for market-wide index volatility, as they are likely to be relevant for estimating the distribution of individual stock returns, which includes both systematic and idiosyncratic volatility components.
3.1.2 Validation
We use the initial period from 1973 to 1994 for hyperparameter optimization. Models with varying hyperparameters are trained from 1973 to 1989 and validated from 1990 to 1994. The hyperparameters yielding the best validation performance, measured in terms of average quantile loss (see Section 3.3 for details), are then applied to train models from 1995 to 2018, generating true out-of-sample predictions. Detailed hyperparameter searches for both full and liquid samples are documented in Appendix D together with the validation scheme.
From 1995 to 2018, using only the U.S. sample, we retrain the model annually using an expanding data window starting in 1973. To prevent data leakage into the next year, the latest date for feature calculation each year is November 29. Separate models are trained for full and liquid samples, with monthly out-of-sample predictions for both U.S. and international samples, using the most recently trained model at each month’s end. Predictions from the full sample model are applied to the full sample and, similarly, to the liquid sample.
3.2 Description of the estimated moments
Since our main empirical results are based on the predicted quantiles and moments of the distribution of all available stocks, we first describe the estimated moments for the US data. As the rest of the international regions show similar dynamics, we present these results in Appendix Table C1. Specifically, we predict 37 $\tau$ quantiles $\tau\in\{0.00005,0.0001,0.001,0.005$, $0.01,\ldots,0.05$, $0.075$, $0.1,0.15,0.2,\ldots,0.8,0.85,0.9$,$0.925$, $0.95,\ldots 0.99$,0.995,0.999,0.9999,0.99995}0.995,0.999,0.9999,0.99995\}0.995 , 0.999 , 0.9999 , 0.99995 }, mean, standard deviation, skewness and kurtosis of all available individual stock returns.
Table 2 contains summary statistics for the predicted moments, and Figure 4 shows the distribution of the predicted mean, standard deviation, skewness and kurtosis for the full and liquid samples for all stocks across the US.
Table 2: Summary statistics for the moments in the U.S. This table shows summary statistics for the forecasted first four moments for the full and liquid samples in the U.S. The moments are forecasted using the quantiles-to-moments algorithm introduced in Section 2.3.
Full Sample Liquid Sample
Mean Median Std. Dev. Mean Median Std. Dev.
Mean 0.0068 0.0089 0.0216 0.0098 0.0092 0.0126
Std. Dev. 0.1477 0.1317 0.0743 0.1071 0.0940 0.0539
Skewness 0.8453 0.7572 0.6352 0.1800 0.1722 0.1993
Kurtosis 8.9784 7.5963 5.2384 5.4261 5.2248 2.3752
We can see that the distribution of the mean for both samples is approximately centred around zero, with the liquid sample having a slightly higher density around the mean. The mean return is positive for both samples, with the liquid sample having a slightly higher mean (0.0098) than the total sample (0.0068). The standard deviation is higher for the total sample (0.1477) than for the liquid sample (0.1071) and shows a wider spread for the total sample than for the liquid sample. Skewness is significantly higher in the full sample (0.8453) than in the liquid sample (0.1800), reflecting more extreme positive and negative returns, while the liquid sample shows a more symmetrical distribution. Kurtosis is also higher in the full sample (8.9784) than in the liquid sample (5.4261), indicating more pronounced tails compared to the liquid sample.
Figure 4: Histogram of Central Distribution Moments. This figure shows histograms of the forecasted first four moments for the full and liquid samples in the U.S. The Moments are forecasted using the quantiles-to-moments algorithm introduced in Section 2.3. Displayed variables are trimmed to exclude the top and bottom 0.5% of their values.
Table 3 also includes average cross-sectional correlations between the first four predicted moments and the median of stock returns. There is a high correlation of $0.95$ between the forecasted mean and the median for the liquid sample, while the correlation is lower ($0.75$) for the full sample. This is consistent with the fact that the full sample contains more micro-cap stocks with higher skewness, which affects the mean but not the median. There is also a high correlation between skewness and kurtosis ($0.91$) in the full sample. In contrast, in the liquid sample, the correlation between skewness and kurtosis is close to zero.
Table 3: Average cross-sectional correlation between variables in the U.S. This table shows the average cross-sectional Spearman correlation between the forecasted first four moments and the median of the stock returns for the full and liquid samples in the U.S. Medians are forecasted using the two-stage model defined in Section 2 and moments are forecasted using the quantiles-to-density algorithm introduced in Section 2.3.
Full Sample Liquid Sample
Median Mean Variance Skewness Kurtosis Median Mean Variance Skewness Kurtosis
Median 1.00 0.75 -0.63 -0.54 -0.44 1.00 0.95 -0.34 0.07 0.16
Mean 0.75 1.00 -0.22 -0.08 -0.08 0.95 1.00 -0.15 0.26 0.09
Variance -0.63 -0.22 1.00 0.53 0.24 -0.34 -0.15 1.00 0.66 -0.30
Skewness -0.54 -0.08 0.53 1.00 0.91 0.07 0.26 0.66 1.00 -0.05
Kurtosis -0.44 -0.08 0.24 0.91 1.00 0.16 0.09 -0.30 -0.05 1.00
3.3 Predictability
Here we evaluate the out-of-sample predictions of the quantiles from our two-stage quantile neural network. We also evaluate the predictions of the mean and variance estimated from the quantiles. This is of particular interest as we argue that the first and second moments, estimated without any assumptions about the distribution, are more predictive of the mean and variance. To benchmark our forecasting strategy, we use a linear model (LNN), the single hidden layer neural network with 32 neurons (1hNN), the double hidden layer neural network with 128 neurons (2hNN) and the commonly used GARCH. Details of the benchmark models can be found in Appendix B.1 and Appendix B.2.
To measure predictive performance across the distribution, we report the average loss, which is defined as the time series average of the cross-sectional $\tau$-quantile losses averaged across all $\tau$s for each month.
$L_{avg}=\frac{1}{T}\sum_{t=1}^{T}\left(\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\left(\frac{1}{K}\sum_{\tau\in\mathcal{T}}L_{\tau,t,i}\right)\right).$ (5)
Table 4 shows the average out-of-sample $\tau$-quantile losses for the full and liquid samples across all regions. The average losses of the two-stage quantile neural network are significantly smaller than all benchmark models when looking at the full sample data. Our two-stage model also significantly outperforms the 1hNN full sample model in all regions except Europe and the 2hNN full sample model in the US and Asia Pacific. The average quantile loss is also lower in Japan, but the difference is not statistically significant at the 5% level.
Predictive performance differences for the liquid sample mirror those observed for the full sample, although to a lesser extent and with fewer statistically significant differences. As expected, average out-of-sample quantile losses are generally lower for the liquid sample than for the full sample, reflecting the lower volatilities of the underlying stocks. The two-stage model consistently outperforms the GARCH model across all regions. However, when compared to alternative neural network specifications, the two-stage model only shows significant outperformance in the US region. Internationally, while the two-stage model achieves lower average quantile losses than the alternative neural network specifications, these differences do not reach statistical significance at the 5% level, except for the 1hNN model in Europe.
Table 4: Out-of-sample quantile forecasts evaluation. This table shows the average out-of-sample quantile cross-section loss for quantile forecasts from the two-stage model and alternative models. The results are obtained separately for the full and liquid samples for individual regions covering the out-of-sample period from 1995 to 2018. Neural network models are trained annually using the expanding window on US data only, while out-of-sample forecasts are generated for both the US and international samples. The average loss is calculated as the time series average of the cross-sectional averages of the average quantile loss across all $\tau\in\{0.00005,0.0001,0.001,0.005$, $0.01,\dots,0.05$, $0.075$, $0.1,0.15,0.2,\dots,0.8,0.85,0.9$,$0.925$, $0.95,\dots 0.99$,0.995,0.999,0.9999,0.99995}0.995,0.999,0.9999,0.99995\}0.995 , 0.999 , 0.9999 , 0.99995 } as in (Eq. 5) for the two-stage NN model, as well as benchmark neural network models (the linear model (LNN), the one-layer model (1hNN), the two-layer model (2hNN)) and the benchmark GARCH model. Note that losses are multiplied by 100.0, and the t-statistics in parentheses are computed for the average loss difference between the given model specification and the two-stage benchmark model using Newey-West standard errors with 12 lags.
Specification Variable Full Sample Liquid Sample
USA Europe Japan Asia Pacific USA Europe Japan Asia Pacific
GARCH $L_{avg}$ 2.659 2.669 2.050 3.416 1.899 1.679 1.836 2.159
(t-stat) (8.47) (14.07) (8.15) (15.51) (6.38) (4.49) (3.05) (6.23)
LNN $L_{avg}$ 2.592 2.520 2.020 3.143 1.874 1.658 1.820 2.115
(t-stat) (6.05) (3.44) (4.43) (4.44) (3.95) (2.07) (1.72) (1.53)
1hNN $L_{avg}$ 2.569 2.496 2.004 3.120 1.869 1.655 1.819 2.121
(t-stat) 3.81 (0.44) (4.03) (3.09) (3.08) (1.69) (1.63) (1.57)
2hNN $L_{avg}$ 2.564 2.494 2.001 3.123 1.870 1.654 1.823 2.128
(t-stat) (2.22) (0.04) (1.71) (2.50) (2.39) (1.10) (1.88) (1.73)
Two stage NN $L_{avg}$ 2.550 2.494 1.992 3.106 1.858 1.649 1.813 2.106
We also examine the performance of the mean and variance forecasts. To compare our two-stage model and the forecasts obtained by the algorithms described earlier, we compute the mean forecasts directly using a two-layer hidden neural network model estimated with a mean square error loss function. We follow Gu et al. (2020) and modify the calculation of the Diebold-Mariano test statistic to compare the cross-sectional average of forecast errors instead of comparing individual errors.
Table 5 shows the out-of-sample $R^{2}$ of the mean and median forecasts. Our two-stage model provides the highest out-of-sample $R^{2}$. For the liquid sample and across all regions, the highest $R^{2}$ is consistently achieved by the median forecast from the two-stage model. Forecasts using the median are not statistically different from those using the mean in the two-stage model. $R^{2}$ for the direct mean prediction from the 2 hidden layers neural network is negative and smaller than the two-stage predictions. In the case of the full sample, the highest $R^{2}$ is consistently achieved by the mean prediction derived from our model.
Table 5: Mean forecast evaluation. This table shows the $R^{2}$ of the mean and median forecasts of the next 22 days of stock returns for the full and liquid samples across all regions. The mean prediction based on the two-stage quantile model (Two-stage NN) is compared with the direct mean prediction from the two-hidden-layer neural network (2hNN MSE - Mean) and the median prediction from the two-stage model (Two-stage NN - Median). Results are obtained separately for full and liquid samples on individual regions covering the period from 1995 to 2018. Models are retrained annually using the expanding window for US data only, while out-of-sample forecasts are generated for both US and international data. Note that losses are multiplied by 100, and values in parentheses are Diebold-Mariano test statistics for the difference between the model in a row and the two-stage NN mean. Standard errors are adjusted using the Newey-West standard deviation estimator with 12 lags.
Specification Full Sample Liquid Sample
USA Europe Japan Asia Pacific Global USA Europe Japan Asia Pacific Global
2hNN MSE - Mean 0.65 0.14 0.85 0.69 0.51 -0.68 -0.44 -0.48 -0.77 -0.54
(-2.54) (-3.56) (-2.37) (-3.39) (-3.55) (-1.81) (-1.31) (-1.84) (-1.79) (-1.98)
Two stage NN - Median 0.10 -0.04 0.30 0.06 0.09 0.37 0.32 0.47 0.55 0.50
(-3.50) (-4.92) (-2.82) (-4.83) (-4.93) (0.23) (0.27) (0.98) (0.39) (0.52)
Two stage NN - Mean 1.63 0.87 1.66 1.63 1.36 0.32 0.22 0.18 0.40 0.37
Finally, to evaluate the second moment forecasts, we use the Mean Absolute Deviation (MAD) and the Root Mean Squared Error (RMSE) of the predicted 22-day volatility for the full and liquid samples for all regions. Table 6 shows a highly significant outperformance of the volatility forecast based on the two-stage model in terms of both MAD and RMSE for all regions and for both the full and liquid samples.
Table 6: Volatility forecast evaluation. This table shows the mean absolute deviation (MAD), root mean squared error (RMSE), and Diebold Mariano test statistic (DM) comparing the corresponding losses of the two competing models for the 22-day volatility forecasts. Volatility forecasting based on two-stage quantile forecasting (two-stage NN) is compared with GARCH model forecasting. Results are obtained separately for full and liquid samples on individual regions covering the period from 1995 to 2018. Neural network models are retrained annually using an expanding window on US data only, while out-of-sample forecasts are generated for both US and international samples. Diebold-Mariano test statistics are computed for the difference between the GARCH and two-stage NN model volatility forecasts. Note that losses are multiplied by 100.0 and the standard errors used in the Diebold-Mariano test statistics are adjusted using the Newey-West standard deviation estimator with 12 lags.
Specification Variable Full Sample Liquid Sample
USA Europe Japan Asia Pacific Global USA Europe Japan Asia Pacific Global
GARCH MAD 7.34 8.79 7.79 11.1 8.53 3.89 3.54 3.80 4.87 3.93
Two stage NN MAD 4.63 5.20 3.99 6.49 4.99 2.93 2.63 2.81 3.16 2.87
GARCH RMSE 13.14 16.24 14.63 18.70 15.50 6.93 7.10 6.72 9.52 7.36
Two stage NN RMSE 7.46 8.52 5.64 10.05 8.08 4.62 4.03 3.98 4.86 4.38
DM (10.25) (35.92) (20.32) (44.26) (26.47) (7.90) (5.98) (7.97) (9.40) (10.02)
Overall, we demonstrate the superior distributional forecasting performance of the two-stage model and the moments derived from it compared to alternative models. The average out-of-sample quantile loss represents the forecast over the entire distribution, while the evaluation of the mean and variance forecasts provides additional evidence derived from the forecasted quantiles.
3.4 Distribution and the cross-section of stock returns
In addition to forecasting, we aim to investigate whether individual distributional characteristics are priced into the cross-section of returns through the profitability of long-short portfolios formed based on forecasted $\tau$-quantiles as well as the profitability of long-short portfolios formed using central moments – mean, volatility, skewness and kurtosis.
3.4.1 $\tau$-quantiles and the cross-section of stock returns
Table 7 shows the average monthly returns and annualised Sharpe ratios for decile long-short equal-weighted portfolios constructed using the $\tau$-quantile forecasts of the two-stage model. Each month, the long and short legs of the portfolios are constructed by selecting the top and bottom 10% of stocks with the highest and lowest forecasts, respectively.
Table 7: Quantiles and the cross-section of stock returns in the U.S. This table shows the average monthly returns and annualised Sharpe ratios for decile long, short and long-short equal-weighted portfolios over the period 1995 to 2018. For each $\tau$-quantile forecast, the long portfolio is formed by taking the top 10% of stocks with the highest forecast each month, The short portfolio is formed by taking the bottom 10% of stocks with the lowest forecast each month. The long-short portfolio is created by taking the difference between the long and short portfolios. Values in brackets are t-statistics adjusted for Newey-West standard errors with 12 lags. Results are obtained separately for the full and liquid samples for the US only.
Full Sample Liquid Sample
Long Short Long-short Long Short Long-short
$\tau$ Mean SR Mean SR Mean SR Mean SR Mean SR Mean SR
0.01 1.22 (5.62) 1.67 0.14 (0.19) 0.04 1.09 (1.52) 0.31 0.95 (5.26) 1.10 0.16 (0.21) 0.05 0.79 (1.06) 0.24
0.05 1.25 (5.66) 1.76 -0.09 (-0.12) -0.02 1.34 (1.86) 0.39 0.96 (5.63) 1.15 0.15 (0.20) 0.04 0.81 (1.11) 0.25
0.1 1.28 (5.73) 1.79 -0.26 (-0.35) -0.07 1.53 (2.13) 0.45 0.98 (5.72) 1.17 0.11 (0.14) 0.03 0.87 (1.20) 0.26
0.2 1.34 (6.04) 1.87 -0.57 (-0.78) -0.16 1.91 (2.68) 0.57 1.02 (5.89) 1.21 0.04 (0.06) 0.01 0.98 (1.34) 0.30
0.3 1.46 (6.54) 1.96 -0.91 (-1.27) -0.27 2.37 (3.36) 0.74 1.10 (6.09) 1.25 -0.04 (-0.06) -0.01 1.14 (1.58) 0.36
0.4 1.88 (7.44) 2.06 -1.34 (-1.91) -0.42 3.22 (4.89) 1.12 1.39 (6.31) 1.34 -0.14 (-0.20) -0.04 1.53 (2.23) 0.54
0.5 2.88 (6.97) 1.78 -1.82 (-2.68) -0.62 4.70 (8.25) 2.27 1.86 (5.61) 1.10 -0.35 (-0.53) -0.12 2.21 (4.21) 1.21
0.6 3.44 (5.68) 1.37 -2.12 (-3.51) -0.88 5.56 (9.41) 3.96 1.74 (3.69) 0.71 -0.38 (-0.88) -0.21 2.11 (5.40) 1.34
0.7 3.48 (4.73) 1.08 -1.42 (-3.57) -1.10 4.90 (6.57) 1.99 1.66 (2.86) 0.56 0.48 (2.50) 0.54 1.18 (2.23) 0.43
0.8 2.98 (3.72) 0.82 0.44 (1.65) 0.62 2.55 (3.16) 0.75 1.28 (1.94) 0.39 0.71 (3.95) 0.85 0.57 (0.91) 0.18
0.9 2.12 (2.57) 0.55 0.75 (3.29) 1.10 1.37 (1.69) 0.38 0.91 (1.24) 0.27 0.76 (4.29) 0.93 0.15 (0.21) 0.05
0.95 1.71 (2.11) 0.45 0.86 (4.04) 1.22 0.85 (1.07) 0.23 0.77 (1.03) 0.23 0.80 (4.62) 0.98 -0.03 (-0.04) -0.01
0.99 1.36 (1.70) 0.36 0.97 (5.04) 1.22 0.39 (0.50) 0.11 0.60 (0.80) 0.17 0.84 (4.72) 1.01 -0.23 (-0.32) -0.07
We show the highest average monthly returns and Sharpe ratios for the long-short portfolios formed based on the central $\tau$-quantiles, $\tau$s of $50$% or $60$%, for both the full sample and the liquid sample. Average monthly returns decrease as we move towards the tails of the distribution. For the full sample, average monthly returns are not significantly different from zero for $\tau$s greater than 80% and less than 10%. For the liquid sample, the average monthly returns fall even more sharply as one moves towards the tails of the distribution, becoming not significantly different from zero for $\tau$s greater than 70% and less than 40%. The highest Sharpe ratios are achieved by the long-short portfolios with the highest average monthly returns.
These results suggest that the tail $\tau$-quantiles in isolation are not priced into the cross-section of stock returns. On the contrary, the central $\tau$-quantiles are associated with the highest average returns. We also provide international evidence showing similar results in Table C2. To further investigate the role of the distribution in pricing the cross-section of stock returns, we turn to the moments of the distribution in the next section.
3.4.2 Distributional moments and the cross-section of stock returns
We then examine the profitability of long-short portfolios formed on the basis of the central moments of the distribution - mean, variance, skewness and kurtosis - forecasts estimated from our two-stage quantile model. To illustrate the importance of asymmetry in the data, we first compare the profitability of portfolios formed on the mean and median forecasts. Table 8 shows the average monthly returns and Sharpe ratios of the long-short decile portfolios based on the mean and median forecasts across all regions.
Table 8: Mean and median portfolios. This table shows average monthly returns and annualised Sharpe ratios for decile long-short equal-weighted portfolios based on the median and mean predictions from the two-stage quantile NN model, the mean prediction from the two-hidden-layer neural network (2hNN - Mean) and the median prediction from the two-stage NN model (2hNN - Median). The results are shown for the full and liquid samples and for individual regions, covering the period from 1995 to 2018. The long portfolio is formed by taking the top 10% of stocks with the highest forecast each month, and the short portfolio is formed by taking the bottom 10% of stocks with the lowest forecast each month. Values in parentheses are t-statistics adjusted for Newey-West standard errors with 12 lags. The significance of the difference in mean returns is tested using t-statistics with Newey-West standard errors with 12 lags.
Specification Variable Full Sample Liquid Sample
USA Europe Japan Asia Pacific Global USA Europe Japan Asia Pacific Global
2hNN MSE - Mean Avg Ret 5.00 3.28 3.11 5.71 4.27 1.62 1.71 1.67 2.05 1.76
SR 3.47 2.87 2.52 3.72 5.50 0.87 1.63 1.52 1.38 2.01
t-stat 4.78 9.26 5.78 5.79 9.93 1.87 1.53 2.81 2.30 2.54
Two stage NN - Median Avg Ret 4.70 3.35 3.10 5.07 4.05 2.21 2.00 2.25 2.95 2.35
SR 2.27 2.48 1.84 3.06 3.53 1.21 1.52 1.51 1.82 2.03
t-stat 4.26 7.14 5.62 8.75 9.16 0.84 0.19 0.10 1.07 0.75
Two stage NN - Mean Avg Ret 6.44 5.29 4.03 8.05 5.95 2.35 2.02 2.27 3.12 2.44
SR 4.34 4.49 3.31 4.99 6.92 1.70 1.88 1.92 2.13 2.83
To benchmark our forecasts, we use the mean forecasts obtained directly from the feed-forward neural network model estimated with the mean squared error loss function. The long-short portfolio based on the mean forecast of our two-stage quantile model leads to significantly higher average monthly returns and Sharpe ratios compared to the portfolios based on the mean forecasts obtained directly from the simple neural network for the full sample across all regions. This is consistent with the $R^{2}$ results in Table 5, as our two-stage model outperforms the direct mean forecast in terms of $R^{2}$ across all regions and also leads to higher corresponding profitability. In the case of the full sample, the average monthly returns of our mean forecast are economically higher than the direct mean forecast in all regions and significantly higher for Japan, Asia-Pacific and globally.
Next, the average monthly returns of the mean- and median-based long-short portfolios are statistically indistinguishable for the liquid sample. This is also consistent with the $R^{2}$ results in Table 5, as there are no significant differences in the quality of the mean and median forecasts for the liquid sample. However, the results are different for the full sample. The difference between the mean and the median is associated with distributional asymmetry, which is more pronounced for stocks in the full sample. Mean-based long-short portfolios have significantly higher average monthly returns than median-based long-short portfolios. This is true for all regions and also globally. The difference between mean and median pricing for the full sample indicates the effect of distributional asymmetry in the cross-section of stock returns.
Table 9: Moments decile portfolios. This table shows average monthly returns for individual deciles (1-10) and long-short equal-weighted portfolios built at the predicted moments. Results are reported for the full and liquid samples, as well as for individual regions, covering the period from 1995 to 2018. Values in parentheses are t-statistics adjusted for Newey-West standard errors with 12 lags.
Mean Full Sample Liquid Sample
Decile USA Europe Japan Asia Pacific USA Europe Japan Asia Pacific
1 -2.37 (-3.87) -1.06 (-1.87) -1.24 (-1.99) -1.69 (-2.40) -0.48 (-0.81) -0.52 (-0.90) -1.13 (-1.99) -1.23 (-1.73)
2 -0.46 (-0.97) -0.29 (-0.63) -0.40 (-0.77) -0.31 (-0.53) 0.33 (0.80) 0.18 (0.44) -0.32 (-0.65) -0.00 (-0.00)
3 0.33 (0.90) 0.04 (0.10) -0.19 (-0.41) 0.13 (0.23) 0.64 (2.00) 0.51 (1.32) -0.01 (-0.02) 0.48 (0.97)
4 0.68 (2.07) 0.35 (0.93) 0.10 (0.24) 0.46 (0.87) 0.75 (2.56) 0.71 (1.97) 0.16 (0.38) 0.58 (1.25)
5 1.01 (3.39) 0.57 (1.64) 0.38 (0.94) 0.87 (1.69) 0.90 (3.21) 0.72 (2.03) 0.20 (0.50) 0.68 (1.53)
6 1.28 (4.37) 0.92 (2.67) 0.70 (1.82) 1.30 (2.57) 1.00 (3.77) 0.91 (2.70) 0.34 (0.96) 1.04 (2.59)
7 1.46 (4.94) 1.22 (3.38) 0.88 (2.27) 1.68 (3.27) 1.17 (4.17) 0.94 (2.57) 0.47 (1.27) 1.10 (2.61)
8 1.82 (5.50) 1.51 (4.05) 1.27 (3.05) 2.30 (4.30) 1.20 (3.77) 1.13 (3.06) 0.65 (1.74) 1.26 (2.83)
9 2.30 (5.92) 2.05 (5.18) 1.63 (3.68) 3.05 (5.02) 1.44 (4.49) 1.37 (3.50) 0.91 (2.45) 1.52 (3.18)
10 4.07 (6.72) 4.23 (8.55) 2.79 (5.44) 6.36 (8.66) 1.87 (5.03) 1.50 (3.55) 1.14 (2.89) 1.88 (3.68)
10-1 6.44 (9.76) 5.29 (12.36) 4.03 (10.26) 8.05 (18.56) 2.35 (5.16) 2.02 (5.51) 2.27 (6.90) 3.11 (7.89)
Median
1 -1.82 (-2.68) -0.42 (-0.62) -0.96 (-1.40) -0.99 (-1.21) -0.35 (-0.54) -0.51 (-0.80) -1.10 (-1.76) -1.09 (-1.43)
2 0.16 (0.29) 0.52 (1.00) 0.08 (0.13) 1.01 (1.42) 0.34 (0.73) 0.23 (0.51) -0.33 (-0.63) -0.10 (-0.18)
3 0.65 (1.48) 0.29 (0.66) 0.25 (0.53) 1.28 (2.01) 0.61 (1.75) 0.50 (1.28) -0.03 (-0.07) 0.50 (1.01)
4 0.99 (2.68) 0.50 (1.31) 0.28 (0.64) 1.14 (1.99) 0.75 (2.44) 0.69 (1.84) 0.17 (0.40) 0.48 (1.00)
5 1.19 (3.74) 0.69 (1.99) 0.36 (0.92) 1.14 (2.23) 0.83 (2.93) 0.77 (2.19) 0.29 (0.78) 0.65 (1.46)
6 1.29 (4.45) 0.91 (2.82) 0.53 (1.44) 1.08 (2.24) 1.08 (3.91) 0.91 (2.67) 0.22 (0.66) 0.93 (2.20)
7 1.44 (5.51) 1.09 (3.38) 0.82 (2.18) 1.37 (2.95) 1.13 (4.29) 0.95 (2.76) 0.46 (1.25) 1.16 (2.81)
8 1.52 (5.26) 1.36 (3.92) 1.04 (2.72) 1.82 (3.83) 1.18 (4.07) 1.03 (2.83) 0.65 (1.87) 1.39 (3.31)
9 1.80 (6.00) 1.65 (4.58) 1.34 (3.54) 2.22 (4.40) 1.41 (4.81) 1.39 (3.80) 0.92 (2.61) 1.53 (3.53)
10 2.88 (6.97) 2.94 (6.73) 2.14 (4.62) 4.08 (6.56) 1.86 (5.61) 1.49 (3.73) 1.16 (3.17) 1.86 (3.81)
10-1 4.70 (8.25) 3.35 (6.64) 3.10 (6.76) 5.07 (10.76) 2.21 (4.22) 2.00 (4.65) 2.25 (5.88) 2.96 (6.27)
Volatility
1 1.02 (4.96) 0.75 (3.23) 0.25 (0.98) 0.76 (2.50) 0.88 (5.13) 0.81 (3.54) 0.33 (1.57) 1.00 (3.14)
2 1.11 (4.28) 0.83 (2.87) 0.36 (1.18) 1.11 (2.89) 0.98 (4.19) 0.84 (2.88) 0.26 (0.90) 0.95 (2.80)
3 1.08 (3.97) 0.92 (2.85) 0.49 (1.44) 1.07 (2.40) 1.02 (4.02) 0.85 (2.84) 0.32 (0.95) 0.91 (2.22)
4 1.09 (3.82) 0.85 (2.38) 0.59 (1.55) 0.98 (1.93) 1.10 (4.09) 0.82 (2.50) 0.27 (0.80) 0.97 (2.20)
5 1.05 (3.24) 0.85 (2.30) 0.60 (1.44) 0.95 (1.73) 0.98 (3.49) 0.81 (2.16) 0.50 (1.38) 0.97 (2.16)
6 1.05 (2.89) 0.66 (1.65) 0.62 (1.39) 0.96 (1.70) 0.96 (3.25) 0.82 (2.11) 0.36 (0.95) 0.79 (1.52)
7 1.05 (2.44) 0.57 (1.25) 0.65 (1.31) 1.08 (1.72) 0.91 (2.41) 0.74 (1.79) 0.26 (0.56) 0.62 (1.24)
8 0.86 (1.65) 0.37 (0.73) 0.83 (1.48) 1.00 (1.37) 0.84 (1.89) 0.84 (1.82) 0.31 (0.64) 0.71 (1.24)
9 0.74 (1.16) 0.62 (1.04) 0.89 (1.41) 1.61 (1.94) 0.72 (1.29) 0.71 (1.33) 0.13 (0.22) 0.62 (0.90)
10 1.05 (1.33) 3.12 (4.21) 0.61 (0.85) 4.63 (4.99) 0.44 (0.58) 0.23 (0.32) -0.34 (-0.45) -0.24 (-0.31)
10-1 0.03 (0.04) 2.36 (3.35) 0.35 (0.62) 3.86 (5.10) -0.44 (-0.60) -0.57 (-0.96) -0.68 (-1.05) -1.24 (-1.81)
Skewness
1 0.85 (3.00) 0.73 (2.07) 0.31 (0.96) 0.69 (1.53) 0.61 (2.65) 0.30 (0.90) -0.06 (-0.19) 0.44 (1.19)
2 1.08 (4.01) 0.88 (2.27) 0.37 (1.04) 0.84 (1.76) 0.75 (2.82) 0.45 (1.20) -0.04 (-0.11) 0.46 (1.07)
3 1.05 (3.50) 0.83 (1.99) 0.48 (1.12) 0.77 (1.50) 0.77 (2.74) 0.58 (1.66) 0.06 (0.14) 0.59 (1.38)
4 1.03 (3.13) 0.61 (1.45) 0.50 (1.08) 0.67 (1.20) 0.81 (2.93) 0.79 (2.26) 0.19 (0.48) 0.70 (1.55)
5 1.04 (3.04) 0.51 (1.21) 0.55 (1.13) 0.97 (1.70) 0.84 (2.94) 0.78 (2.18) 0.20 (0.48) 0.86 (1.85)
6 0.88 (2.51) 0.59 (1.50) 0.56 (1.12) 1.31 (2.26) 1.02 (3.23) 0.84 (2.47) 0.42 (1.03) 0.74 (1.52)
7 0.90 (2.27) 0.71 (1.74) 0.67 (1.38) 1.67 (2.70) 1.11 (3.45) 1.00 (2.79) 0.38 (0.85) 1.02 (2.15)
8 0.96 (2.18) 0.99 (2.32) 0.87 (1.74) 2.20 (3.02) 1.09 (2.95) 0.97 (2.35) 0.52 (1.20) 1.01 (2.05)
9 1.23 (2.45) 1.65 (3.83) 0.98 (1.92) 2.67 (3.65) 1.03 (2.34) 1.04 (2.26) 0.54 (1.25) 0.95 (1.54)
10 1.09 (1.92) 2.03 (4.92) 0.61 (1.30) 2.37 (3.63) 0.80 (1.10) 0.72 (1.16) 0.19 (0.35) 0.54 (0.75)
10-1 0.24 (0.57) 1.30 (4.07) 0.30 (1.08) 1.68 (4.06) 0.18 (0.28) 0.42 (1.02) 0.25 (0.72) 0.11 (0.20)
Kurtosis
1 1.08 (2.69) 0.82 (1.90) 0.48 (1.01) 0.80 (1.46) 0.76 (1.26) 0.39 (0.61) 0.03 (0.04) 0.47 (0.70)
2 0.99 (3.32) 0.93 (2.28) 0.47 (1.26) 0.68 (1.34) 0.70 (1.44) 0.67 (1.34) -0.17 (-0.32) 0.45 (0.70)
3 1.00 (3.35) 0.76 (1.76) 0.51 (1.25) 0.69 (1.28) 0.70 (1.85) 0.63 (1.44) 0.11 (0.24) 0.34 (0.61)
4 1.05 (3.40) 0.69 (1.62) 0.49 (1.11) 0.90 (1.58) 0.80 (2.43) 0.73 (1.98) 0.23 (0.57) 0.72 (1.57)
5 0.96 (2.76) 0.54 (1.29) 0.57 (1.19) 1.07 (1.93) 0.78 (2.49) 0.90 (2.48) 0.37 (0.99) 0.76 (1.68)
6 0.99 (2.64) 0.72 (1.68) 0.70 (1.37) 1.78 (2.83) 0.91 (3.19) 0.78 (2.27) 0.43 (1.32) 0.99 (2.34)
7 0.99 (2.27) 0.84 (2.05) 0.74 (1.51) 1.94 (2.91) 1.13 (4.45) 0.90 (2.71) 0.50 (1.46) 1.00 (2.30)
8 1.02 (2.30) 1.31 (3.09) 0.91 (1.82) 2.36 (3.28) 1.07 (3.98) 0.90 (2.67) 0.37 (1.10) 0.88 (2.12)
9 1.19 (2.42) 1.49 (4.03) 0.76 (1.62) 2.41 (3.68) 1.07 (3.77) 0.94 (2.85) 0.37 (1.07) 1.02 (2.29)
10 0.83 (2.16) 1.42 (4.54) 0.27 (0.64) 1.52 (2.95) 0.92 (3.37) 0.62 (1.90) 0.16 (0.47) 0.69 (1.63)
10-1 -0.25 (-0.77) 0.60 (1.82) -0.21 (-0.66) 0.73 (1.80) 0.15 (0.33) 0.23 (0.62) 0.14 (0.33) 0.22 (0.57)
In addition to examining the role of individual $\tau$-quantiles and the mean, we also look at the decile portfolios and the corresponding long-short portfolios formed on the basis of the higher moments. The higher moments are obtained by the quantiles-to-moments algorithm using the two-stage model quantile forecasts as described in Section 2.3. Table 9 shows the average monthly returns and Sharpe ratios of the equally weighted decile long-short portfolios based on the forecast mean, median, volatility, skewness and kurtosis.
Overall, none of the higher moments appear to be monotonically priced in the cross-section of stock returns in the liquid or full sample. Average monthly returns of the long-short portfolios formed on the higher moments are mostly not significantly different from zero. There are some regional exceptions for the full sample, where volatility and skewness-based long-short portfolios have significantly positive average monthly returns in Europe and Asia Pacific. Results for the value-weighted long-short portfolios are mostly consistent with the equal-weighted portfolios and are shown in Table C3 in Appendix.
The results on volatility show a negative relationship between predicted volatility and future return for the liquid sample, in line with previous academic research. However, the relationship is not statistically significant, which is to be expected given the focus on large cap stocks. The full sample results are plagued by market microstructure distortions for equal-weighted portfolios, such as bid-ask jumps, and the pattern is clearer for value-weighted portfolios. However, the effect remains statistically insignificant, which can be explained by our sample period, which focuses primarily on years more recent than the sample period of Ang et al. (2006), the original study that identified volatility as priced into the cross-section of returns.
The results for skewness are inconclusive for a monotonic relationship with future returns. However, there is a noticeable hill-shaped effect; stocks with low skewness have lower future returns relative to stocks with more average skewness. The low returns of the high-skewness portfolio (10) lead to a non-monotonic relationship across the decile portfolios and thus also affects the 10-1 long-short portfolio. Skewness would be more convincingly priced if 9-1 long-short portfolios were used instead.
4 Conclusion
We develop a novel machine learning approach to forecasting the full distribution of stock returns, focusing on both US and international markets. Our method uses a multi-head quantile neural network to predict a set of $\tau$ quantiles based on stock characteristics and market variables. We interpolate these quantiles using cubic B-splines to derive the full distribution and its moments, such as mean, volatility, skewness and kurtosis. This approach outperforms traditional parametric models and simpler neural networks in forecasting stock returns, providing more accurate out-of-sample mean and variance forecasts.
We also examine the pricing of distributional characteristics in the cross-section of stock returns using a comprehensive dataset of US and international stocks. We find that only the central $\tau$-quantiles are priced, with significant profitability of long-short portfolios based on these predictions. In addition, our results show that the mean obtained by numerical integration of the interpolated distribution leads to higher out-of-sample returns than direct mean forecasts from other models. We find no evidence that other moments, such as variance, skewness and kurtosis, are priced in the cross-section, adding to the ongoing debate in the asset pricing literature.
Appendix for
“Predicting the distributions of stock returns around the globe in the era of big data and learning”
Appendix A List of the anomalies
Table A.1: List of anomalies
Fundamental
Accruals
Accruals Sloan (1996)
Change in Common Equity Richardson et al. (2006)
Change in Current Operating Assets Richardson et al. (2006)
Change in Current Operating Liabilities Richardson et al. (2006)
Change in Financial Liabilities Richardson et al. (2006)
Change in Long-Term Investments Richardson et al. (2006)
Change in Net Financial Assets Richardson et al. (2006)
Change in Net Non-Cash Working Capital Richardson et al. (2006)
Change in Net Non-Current Operating Assets Richardson et al. (2006)
Change in Non-Current Operating Assets Richardson et al. (2006)
Change in Non-Current Operating Liabilities Richardson et al. (2006)
Change in Short-Term Investments Richardson et al. (2006)
Discretionary Accruals Dechow et al. (1995)
Growth in Inventory Thomas and Zhang (2002)
Inventory Change Thomas and Zhang (2002)
Inventory Growth Belo and Lin (2011)
M/B and Accruals Bartov and Kim (2004)
Net Working Capital Changes Soliman (2008)
Percent Operating Accrual Hafzalla et al. (2011)
Percent Total Accrual Hafzalla et al. (2011)
Total Accruals Richardson et al. (2006)
Intangibles
$\bigtriangleup$ Gross Margin - $\bigtriangleup$ Sales Abarbanell and Bushee (1998)
$\bigtriangleup$ Sales - $\bigtriangleup$ Accounts Receivable Abarbanell and Bushee (1998)
$\bigtriangleup$ Sales - $\bigtriangleup$ Inventory Abarbanell and Bushee (1998)
$\bigtriangleup$ Sales - $\bigtriangleup$ SG and A Abarbanell and Bushee (1998)
Asset Liquidity Ortiz-Molina and Phillips (2014)
Asset Liquidity II Ortiz-Molina and Phillips (2014)
Cash-to-assets Palazzo (2012)
Earnings Conservatism Francis et al. (2004)
Earnings Persistence Francis et al. (2004)
Earnings Predictability Francis et al. (2004)
Earnings Smoothness Francis et al. (2004)
Earnings Timeliness Francis et al. (2004)
Herfindahl Index Hou and Robinson (2006)
Hiring rate Belo et al. (2014)
Industry Concentration Assets Hou and Robinson (2006)
Industry Concentration Book Equity Hou and Robinson (2006)
Industry-adjusted Organizational Capital-to-Assets Eisfeldt and Papanikolaou (2013)
Industry-adjusted Real Estate Ratio Tuzel (2010)
Org. Capital Eisfeldt and Papanikolaou (2013)
RD / Market Equity Chan et al. (2001)
RD Capital-to-assets Li (2011)
RD Expenses-to-sales Chan et al. (2001)
Tangibility Hahn and Lee (2009)
Unexpected RD Increases Eberhart et al. (2004)
Whited-Wu Index Whited and Wu (2006)
Investment
$\bigtriangleup$ CAPEX - $\bigtriangleup$ Industry CAPEX Abarbanell and Bushee (1998)
Asset Growth Cooper et al. (2008)
Change Net Operating Assets Hirshleifer et al. (2004)
Changes in PPE and Inventory-to-Assets Lyandres et al. (2007)
Composite Debt Issuance Lyandres et al. (2007)
Composite Equity Issuance (5-Year) Daniel and Titman (2006)
Debt Issuance Spiess and Affleck-Graves (1995)
Growth in LTNOA Fairfield et al. (2003)
Investment Titman et al. (2004)
Net Debt Finance Bradshaw et al. (2006)
Net Equity Finance Bradshaw et al. (2006)
Net Operating Assets Hirshleifer et al. (2004)
Noncurrent Operating Assets Changes Soliman (2008)
Share Repurchases Ikenberry et al. (1995)
Total XFIN Bradshaw et al. (2006)
Profitability
Asset Turnover Soliman (2008)
Capital Turnover Haugen and Baker (1996)
Cash-based Operating Profitability Ball et al. (2016)
Change in Asset Turnover Soliman (2008)
Change in Profit Margin Soliman (2008)
Earnings / Price Basu (1977)
Earnings Consistency Alwathainani (2009)
F-Score Piotroski (2000)
Gross Profitability Novy-Marx (2013)
Labor Force Efficiency Abarbanell and Bushee (1998)
Leverage Bhandari (1988)
O-Score (More Financial Distress) Dichev (1998)
Operating Profits to Assets Ball et al. (2016)
Operating Profits to Equity Fama and French (2015)
Profit Margin Soliman (2008)
Return on Net Operating Assets Soliman (2008)
Return-on-Equity Haugen and Baker (1996)
Z-Score (Less Financial Distress) Dichev (1998)
Value
Assets-to-Market Fama and French (1992)
Book Equity / Market Equity Fama and French (1992)
Cash Flow / Market Equity Lakonishok et al. (1994)
Duration of Equity Dechow et al. (2004)
Enterprise Component of Book/Price Penman et al. (2007)
Enterprise Multiple Loughran and Wellman (2011)
Intangible Return Daniel and Titman (2006)
Leverage Component of Book/Price Penman et al. (2007)
Net Payout Yield Boudoukh et al. (2007)
Operating Leverage Novy-Marx (2010)
Payout Yield Boudoukh et al. (2007)
Sales Growth Lakonishok et al. (1994)
Sales/Price Barbee Jr et al. (1996)
Sustainable Growth Lockwood and Prombutr (2010)
Market Friction
11-Month Residual Momentum Blitz et al. (2011)
52-Week High George and Hwang (2004)
Amihud’s Measure (Illiquidity) Amihud (2002)
Beta Fama and MacBeth (1973)
Betting against Beta Frazzini and Pedersen (2014)
Bid-Ask Spread Amihud and Mendelson (1986)
Cash Flow Variance Haugen and Baker (1996)
Coefficient of Variation of Share Turnover Chordia et al. (2001)
Coskewness Harvey and Siddique (2000)
Downside Beta Ang et al. (2006)
Earnings Forecast-to-Price Elgers et al. (2001)
Firm Age Barry and Brown (1984)
Firm Age-Momentum Zhang (2006)
Idiosyncratic Risk Ang et al. (2006)
Industry Momentum Moskowitz and Grinblatt (1999)
Lagged Momentum Novy-Marx (2012)
Liquidity Beta 1 Acharya and Pedersen (2005)
Liquidity Beta 2 Acharya and Pedersen (2005)
Liquidity Beta 3 Acharya and Pedersen (2005)
Liquidity Beta 4 Acharya and Pedersen (2005)
Liquidity Beta 5 Acharya and Pedersen (2005)
Liquidity Shocks Bali et al. (2013)
Long-Term Reversal Bondt and Thaler (1985)
Max Bali et al. (2011)
Momentum Jegadeesh and Titman (1993)
Momentum and LT Reversal Kot and Chan (2006)
Momentum-Reversal Jegadeesh and Titman (1993)
Momentum-Volume Lee and Swaminathan (2000)
Price Blume and Husic (1973)
Seasonality Heston and Sadka (2008)
Seasonality 1 A Heston and Sadka (2008)
Seasonality 1 N Heston and Sadka (2008)
Seasonality 11-15 A Heston and Sadka (2008)
Seasonality 11-15 N Heston and Sadka (2008)
Seasonality 16-20 A Heston and Sadka (2008)
Seasonality 16-20 N Heston and Sadka (2008)
Seasonality 2-5 A Heston and Sadka (2008)
Seasonality 2-5 N Heston and Sadka (2008)
Seasonality 6-10 A Heston and Sadka (2008)
Seasonality 6-10 N Heston and Sadka (2008)
Share Issuance (1-Year) Pontiff and Woodgate (2008)
Share Turnover Datar et al. (1998)
Short-Term Reversal Jegadeesh (1990)
Size Banz (1981)
Tail Risk Kelly and Jiang (2014)
Total Volatility Ang et al. (2006)
Volume / Market Value of Equity Haugen and Baker (1996)
Volume Trend Haugen and Baker (1996)
Volume Variance Chordia et al. (2001)
I/B/E/S
Analyst Value Frankel and Lee (1998)
Analysts Coverage Elgers et al. (2001)
Change in Forecast + Accrual Barth and Hutton (2004)
Change in Recommendation Jegadeesh et al. (2004)
Changes in Analyst Earnings Forecasts Hawkins et al. (1984)
Disparity between LT and ST Earnings Growth Forecasts Da and Warachka (2011)
Dispersion in Analyst LT Growth Forecasts Anderson et al. (2005)
Down Forecast Barber et al. (2001)
Forecast Dispersion Diether et al. (2002)
Long-Term Growth Forecasts La Porta (1996)
Up Forecast Barber et al. (2001)
Appendix B Alternative models
We describe the benchmark models used to evaluate the performance of our two-stage model and the accuracy of our distributional moment forecasts. We compare the forecasting performance of the two-stage model with the GARCH model and three different neural network architectures, as outlined in Table 4. Detailed descriptions of the neural networks and the GARCH model are provided in Sections B.1 and B.2, respectively.
B.1 Feed-forward neural networks
We use multi-head feed-forward neural networks with varying architectures as benchmarks for the two-stage model quantile forecasts. We consider a simple multi-output linear model without any hidden layers, neural networks with 1 hidden layer, and neural networks with 2 hidden layers as benchmarks for the two-stage model. Each model has 176 features and 37 outputs. The input features are the same as those for the standardized $\tau$-quantiles network in the two-stage model. The 37 outputs are the $\tau$-quantile forecasts of raw future stock returns for the same $\tau$ values as in the two-stage model. See Section 3.1.1 for the feature specification and Section 2 for the main model introduction.
Each neural network is trained using the aggregated multi-$\tau$-quantile loss.
$\mathcal{L}^{\mathcal{T}}_{raw}=\frac{1}{B}\frac{1}{K}\sum_{\tau\in\mathcal{T}}\sum_{i=1}^{B}\rho_{\tau}\left(r_{i,t}-\widehat{Q}_{r_{i,t}}(\tau)\right)$ (6)
where $\widehat{Q}_{r_{i,t}}(\tau)$ is the forecast $\tau$-quantile of the raw return of stock $i$ at period $t$, $B$ is the lot size and $K=37$ is the number of $\tau$ in the $\mathcal{T}$ set.
We also use a single-head feed-forward neural network with 2 hidden layers and mean squared error loss function as a benchmark for mean forecasts comparison as in Table 5 and Table 8. This model also uses the same 176 features as the networks above. Architecture schemas for all alternative neural network models are shown in Figure B.1.
(176) $Q_{r}(\tau)$ (37)
(a) Linear model
(176) (128) $Q_{r}(\tau)$ (37)
(b) One hidden layer neural network
(176) (128) (128) $Q_{r}(\tau)$ (37)
(c) Two hidden layers neural network
(176) (128) (128) $\mu$ (1)
(d) Single output neural network with two hidden layers
Figure B.1: Architectures of benchmark multi-head feed-forward neural networks.
1(a) depicts the linear model, 1(b) illustrates the neural network with one hidden layer, 1(c) shows the neural network with two hidden layers, and 1(d) presents the two hidden layers neural network with a single output node, which is used to generate mean forecasts directly using the mean squared error loss function. All networks have 176 input features; multi-output networks have 37 outputs, varying in the number of hidden layers and neurons.
Most of the hyper-parameters for all four benchmark neural networks are the same as for the two-stage model and are available in Table D1. Dropout rate is set to 0.2 and learning rate to 0.0003, in line with the selected hyper-parameter for the two-stage model. Hyper-parameters that differ are the architecture in terms of the number of layers and neurons, the loss function, and other model-specific settings.
B.2 GARCH benchmark
In order to provide a valid benchmark for the two-stage model, we examine various specifications of the GARCH model. We utilize daily returns from all regions for the period from 1995 to 2018, corresponding to the time period and stock universe used for the main results in this paper. We base our estimations on daily returns using a rolling window of 36 months. Results are compared for both the full and liquid samples.
We consider several variants of GARCH(1,1) model with the standardized residuals distributed according to normal and t-distributions. Mean $\mu$ is either estimated or set equal to the risk-free rate at the end of the estimation window plus 5% divided by 252. Assuming these distributions, we simulate 100,000 paths of residuals for 22 days ahead and consequently derive the corresponding return paths. During this process, volatility is simulated using Equation 7
$\sigma_{i,j}^{2}=\omega+\alpha\epsilon_{i-1,j}^{2}+\beta\sigma_{i-1,j}^{2}$ (7)
where $\omega$, $\alpha$, and $\beta$ are the GARCH parameters, $\epsilon_{i-1,j}$ represents the previous day’s residual, and $\sigma_{i-1,j}^{2}$ is the previous day’s volatility. The cumulative return for the $j$-th simulated path is then calculated according to Equation 8.
$r_{j}=\left(\prod_{i=1}^{22}\left(1+\mu+\epsilon_{i,j}\sqrt{\sigma_{i,j}^{2}}\right)\right)-1$ (8)
Finally, the variance forecast is taken as the sample standard deviation of all the cumulative returns over the 22 days, and the quantile forecast is obtained as the sample quantile of all the cumulative returns over the 22 days.
Table B1: Performance of Various GARCH Specifications This table shows out-of-sample average quantile losses for different specifications of the GARCH model for both the full and liquid samples. The significance of the difference in average quantile losses between the GARCH(1,1) with Student’s t-distribution and other specifications is captured by t-statistics adjusted for Newey-West standard errors with 12 lags. All models, unless specified, use the mean as a risk-free rate at the end of the estimation window plus 5%. GARCH(1,1) t-dist uses a t-distribution with an estimated number of degrees of freedom, and GARCH(1,1) normal dist uses a normal distribution. GARCH(1,1) fixed, t-dist models have the number of degrees of freedom fixed at 3, 4, or 5, and their auto-regressive parameters are fixed to $\alpha=0.06$ and $\beta=0.94$. Daily risk-free rate data are taken from the Fama-French data library.
Specification Variable Full Sample Liquid Sample
GARCH(1, 1), t-dist Avg Loss 0.02642 0.01887
GARCH(1, 1) fixed, t-dist 3 df Avg Loss 0.02636 0.01906
t-stat -0.96 4.97
GARCH(1, 1) fixed, t-dist 4 df Avg Loss 0.02641 0.01900
t-stat -0.17 4.25
GARCH(1, 1) fixed, t-dist 5 df Avg Loss 0.02646 0.01901
t-stat 1.06 4.53
GARCH(1, 1), normal dist Avg Loss 0.02681 0.01899
t-stat 15.25 10.30
GARCH(1, 1), t-dist, $\mu$ estimated Avg Loss 0.02769 0.01930
t-stat 5.22 4.23
GARCH(2, 2), t-dist Avg Loss 0.02649 0.01893
t-stat 2.10 1.26
EGARCH(1, 1, 1), t-dist Avg Loss 0.02681 0.01913
t-stat 3.38 2.25
For the t-distribution, besides estimating the number of degrees of freedom from the data, we also consider fixing the number of degrees of freedom to 3, 4, and 5. Furthermore, we try GARCH(2, 2) and EGARCH(1, 1, 1) specifications also with t-distributed standardized residuals, in both cases estimating degrees of freedom from the data.
[…]
A more recent method for quantile prediction is MIDAS quantile regression introduced in Ghysels et al. (2016). MIDAS quantile regressions, in their simplest form, predict quantile as $Q=\alpha+\beta f(\gamma)$. $\alpha$, $\beta$, and $\gamma$ are three parameters to be estimated via non-linear OLS. $f(\gamma)$ is a weighted average of past absolute daily returns where the parameter $\gamma$ determines the speed of weight decay for less recent returns. The simplest specification is quite similar to GARCH(1, 1) with its auto-regressive component. In contrast to GARCH, MIDAS quantile regression does not rely on any distributional assumptions, which, however, also means that it cannot be applied to the extreme quantiles being estimated in the present study. The range of quantities that can be estimated via MIDAS is limited by a number of observations for one stock. Three years of data for one stock allows, at most, an estimation of 0.0014 quantile, which is much larger than the minimum quantiles estimated by neural networks 0.00005.
Appendix C Additional empirical results
Figure C.1: The cross-sectional average of volatilities as a scaling factor. This figure shows the evolution over time of the cross-sectional average of stock-level volatilities for the full sample of U.S. stocks. Stock-level volatilities are calculated as exponential moving averages of squared daily returns with a smoothing factor of $0.94$. The scaling factor is used to standardize stock returns to be used as a label for the standardized $\tau$-quantile neural network or stage 1 of the two-stage model defined in Section 2.
Table C1: Average cross-sectional correlation between moments and median internationally. This table presents the average cross-sectional Spearman correlation between the forecasted first four moments and the median of the stock returns for the full and liquid samples in Europe, Japan, and Asia Pacific. Medians are forecasted using the two-stage model defined in Section 2, and moments are forecasted using the quantiles-to-moments algorithm defined in Section 2.3. The predictions span the time period from 1995 to 2018.
Europe
Full Sample Liquid Sample
Median Mean Variance Skewness Kurtosis Median Mean Variance Skewness Kurtosis
Median 1.0 0.81 -0.58 -0.49 -0.43 1.0 0.97 -0.2 0.32 0.18
Mean 0.81 1.0 -0.25 -0.13 -0.16 0.97 1.0 -0.05 0.46 0.11
Variance -0.58 -0.25 1.0 0.43 0.2 -0.2 -0.05 1.0 0.49 -0.47
Skewness -0.49 -0.13 0.43 1.0 0.94 0.32 0.46 0.49 1.0 -0.1
Kurtosis -0.43 -0.16 0.2 0.94 1.0 0.18 0.11 -0.47 -0.1 1.0
Japan
Full Sample Liquid Sample
Median Mean Variance Skewness Kurtosis Median Mean Variance Skewness Kurtosis
Median 1.0 0.83 -0.51 -0.39 -0.32 1.0 0.97 -0.27 0.27 0.23
Mean 0.83 1.0 -0.19 -0.02 -0.04 0.97 1.0 -0.12 0.42 0.17
Variance -0.51 -0.19 1.0 0.31 0.04 -0.27 -0.12 1.0 0.48 -0.44
Skewness -0.39 -0.02 0.31 1.0 0.93 0.27 0.42 0.48 1.0 0.0
Kurtosis -0.32 -0.04 0.04 0.93 1.0 0.23 0.17 -0.44 0.0 1.0
Asia Pacific
Full Sample Liquid Sample
Median Mean Variance Skewness Kurtosis Median Mean Variance Skewness Kurtosis
Median 1.0 0.8 -0.65 -0.54 -0.43 1.0 0.96 -0.24 0.19 0.18
Mean 0.8 1.0 -0.29 -0.18 -0.16 0.96 1.0 -0.07 0.36 0.1
Variance -0.65 -0.29 1.0 0.52 0.26 -0.24 -0.07 1.0 0.61 -0.42
Skewness -0.54 -0.18 0.52 1.0 0.92 0.19 0.36 0.61 1.0 -0.1
Kurtosis -0.43 -0.16 0.26 0.92 1.0 0.18 0.1 -0.42 -0.1 1.0
Table C2: Quantiles and long-short portfolios internationally. This table displays average monthly returns and annualized Sharpe ratios (SR) for decile long, short, and long-short equal-weighted portfolios over the period from 1995 to 2018. For each $\tau$-quantile forecast, the long portfolio is formed by selecting the top 10% of stocks with the highest forecasts each month, the short portfolio by selecting the bottom 10% of stocks with the lowest forecasts, and the long-short portfolio by taking the difference between the long and short portfolios. Values in brackets are t-statistics adjusted for Newey-West standard errors with 12 lags. Results are obtained separately for full and liquid samples for Europe, Japan, and Asia Pacific.
Europe
Full Sample Liquid Sample
Long Short Long-Short Long Short Long-Short
$\tau$ Mean SR Mean SR Mean SR Mean SR Mean SR Mean SR
0.01 0.91 (3.92) 1.04 2.57 (3.61) 1.07 -1.66 (-2.41) -0.77 0.98 (4.06) 0.99 0.06 (0.09) 0.02 0.92 (1.51) 0.39
0.05 0.86 (3.79) 1.00 2.32 (3.22) 0.96 -1.46 (-2.10) -0.66 1.00 (4.31) 1.03 -0.01 (-0.01) -0.00 1.01 (1.63) 0.42
0.1 0.88 (3.89) 1.02 2.13 (2.94) 0.88 -1.25 (-1.80) -0.57 1.03 (4.53) 1.05 -0.05 (-0.07) -0.02 1.09 (1.79) 0.46
0.2 0.95 (4.21) 1.11 1.70 (2.37) 0.71 -0.75 (-1.10) -0.35 1.09 (4.53) 1.09 -0.14 (-0.20) -0.05 1.23 (2.03) 0.55
0.3 1.08 (4.73) 1.23 1.19 (1.67) 0.51 -0.10 (-0.16) -0.05 1.22 (4.76) 1.15 -0.28 (-0.38) -0.10 1.50 (2.54) 0.70
0.4 1.69 (5.95) 1.59 0.35 (0.50) 0.16 1.35 (2.20) 0.74 1.39 (4.76) 1.12 -0.33 (-0.47) -0.13 1.71 (3.12) 0.91
0.5 2.94 (6.73) 1.80 -0.42 (-0.62) -0.20 3.35 (6.63) 2.48 1.49 (3.73) 0.88 -0.50 (-0.80) -0.22 2.00 (4.65) 1.52
0.6 3.65 (6.89) 1.82 -0.67 (-1.17) -0.37 4.32 (10.29) 3.48 1.49 (3.05) 0.70 -0.31 (-0.65) -0.19 1.80 (5.33) 1.47
0.7 3.98 (6.39) 1.70 -0.57 (-1.69) -0.51 4.55 (9.80) 2.55 1.29 (2.21) 0.52 0.23 (0.78) 0.21 1.06 (2.63) 0.59
0.8 4.12 (5.82) 1.63 0.17 (0.74) 0.20 3.95 (6.10) 1.74 0.97 (1.51) 0.36 0.54 (2.23) 0.53 0.43 (0.85) 0.20
0.9 3.70 (5.07) 1.46 0.45 (1.93) 0.54 3.25 (4.70) 1.41 0.70 (1.01) 0.25 0.64 (2.66) 0.64 0.06 (0.11) 0.03
0.95 3.54 (4.78) 1.40 0.57 (2.46) 0.66 2.98 (4.27) 1.31 0.54 (0.76) 0.19 0.67 (2.80) 0.68 -0.13 (-0.23) -0.06
0.99 3.40 (4.69) 1.39 0.75 (2.98) 0.78 2.65 (3.97) 1.22 0.42 (0.58) 0.15 0.70 (2.96) 0.71 -0.28 (-0.48) -0.12
Japan
Full Sample Liquid Sample
Long Short Long-Short Long Short Long-Short
$\tau$ Mean SR Mean SR Mean SR Mean SR Mean SR Mean SR
0.01 0.62 (2.24) 0.61 -0.04 (-0.06) -0.01 0.66 (1.21) 0.25 0.50 (2.44) 0.43 -0.57 (-0.75) -0.19 1.07 (1.68) 0.40
0.05 0.63 (2.22) 0.62 -0.16 (-0.22) -0.05 0.79 (1.45) 0.30 0.53 (2.62) 0.46 -0.60 (-0.78) -0.20 1.13 (1.76) 0.43
0.1 0.67 (2.31) 0.65 -0.25 (-0.35) -0.08 0.92 (1.73) 0.35 0.61 (2.91) 0.53 -0.63 (-0.84) -0.21 1.24 (1.96) 0.47
0.2 0.79 (2.63) 0.76 -0.39 (-0.55) -0.13 1.18 (2.29) 0.46 0.71 (3.32) 0.60 -0.64 (-0.85) -0.22 1.35 (2.16) 0.53
0.3 1.01 (3.27) 0.93 -0.53 (-0.75) -0.18 1.54 (3.13) 0.63 0.88 (3.76) 0.71 -0.77 (-1.04) -0.26 1.65 (2.78) 0.68
0.4 1.49 (4.33) 1.22 -0.72 (-1.03) -0.25 2.21 (4.77) 1.01 1.08 (4.04) 0.76 -0.88 (-1.24) -0.31 1.96 (3.72) 0.93
0.5 2.14 (4.62) 1.29 -0.96 (-1.40) -0.35 3.10 (6.76) 1.84 1.16 (3.17) 0.67 -1.10 (-1.76) -0.44 2.25 (5.87) 1.51
0.6 2.29 (4.09) 1.08 -1.16 (-1.92) -0.50 3.45 (7.98) 2.96 1.05 (2.12) 0.50 -0.98 (-2.07) -0.53 2.03 (6.33) 1.83
0.7 2.26 (3.58) 0.90 -1.03 (-2.25) -0.70 3.28 (6.44) 1.98 0.74 (1.19) 0.30 -0.31 (-0.99) -0.24 1.05 (2.16) 0.55
0.8 1.86 (2.69) 0.66 -0.44 (-1.30) -0.40 2.30 (3.95) 0.98 0.34 (0.48) 0.12 0.02 (0.08) 0.02 0.32 (0.55) 0.13
0.9 1.34 (1.87) 0.44 -0.14 (-0.50) -0.14 1.49 (2.49) 0.57 0.08 (0.11) 0.03 0.16 (0.69) 0.14 -0.08 (-0.14) -0.03
0.95 1.11 (1.57) 0.36 -0.01 (-0.04) -0.01 1.13 (1.94) 0.42 -0.04 (-0.05) -0.01 0.20 (0.89) 0.18 -0.24 (-0.39) -0.09
0.99 0.92 (1.33) 0.30 0.18 (0.73) 0.17 0.75 (1.34) 0.29 -0.15 (-0.20) -0.05 0.24 (1.08) 0.21 -0.39 (-0.61) -0.14
Asia Pacific
Full Sample Liquid Sample
Long Short Long-Short Long Short Long-Short
$\tau$ Mean SR Mean SR Mean SR Mean SR Mean SR Mean SR
0.01 1.01 (3.33) 0.95 3.57 (4.15) 1.11 -2.56 (-3.63) -0.99 1.18 (3.77) 0.79 -0.50 (-0.64) -0.14 1.68 (2.36) 0.54
0.05 1.02 (3.34) 0.98 2.94 (3.42) 0.92 -1.92 (-2.82) -0.75 1.18 (3.74) 0.79 -0.61 (-0.77) -0.17 1.79 (2.51) 0.58
0.1 1.05 (3.42) 1.01 2.47 (2.87) 0.78 -1.41 (-2.11) -0.56 1.23 (3.92) 0.83 -0.67 (-0.84) -0.19 1.90 (2.66) 0.62
0.2 1.10 (3.51) 1.05 1.64 (1.90) 0.53 -0.54 (-0.82) -0.22 1.31 (4.20) 0.87 -0.70 (-0.88) -0.20 2.01 (2.86) 0.68
0.3 1.28 (3.91) 1.16 0.71 (0.84) 0.23 0.57 (0.91) 0.25 1.52 (4.56) 0.96 -0.83 (-1.04) -0.24 2.35 (3.39) 0.84
0.4 2.30 (5.37) 1.55 -0.18 (-0.21) -0.06 2.48 (4.54) 1.24 1.81 (5.39) 1.07 -0.90 (-1.15) -0.27 2.71 (4.65) 1.23
0.5 4.08 (6.56) 1.75 -0.99 (-1.21) -0.35 5.07 (10.75) 3.06 1.86 (3.81) 0.76 -1.09 (-1.43) -0.36 2.95 (6.28) 1.82
0.6 5.31 (6.81) 1.74 -1.38 (-1.82) -0.55 6.69 (14.28) 3.55 1.54 (2.62) 0.51 -0.59 (-1.04) -0.27 2.13 (7.48) 1.10
0.7 6.04 (7.17) 1.75 -1.33 (-2.43) -0.82 7.37 (13.30) 2.84 1.29 (2.05) 0.39 0.43 (1.23) 0.28 0.86 (2.02) 0.31
0.8 6.02 (6.81) 1.67 -0.01 (-0.02) -0.01 6.03 (8.91) 1.97 0.83 (1.21) 0.24 0.71 (2.16) 0.49 0.12 (0.22) 0.04
0.9 5.69 (6.24) 1.58 0.41 (1.30) 0.39 5.29 (7.41) 1.75 0.48 (0.68) 0.13 0.79 (2.46) 0.55 -0.30 (-0.51) -0.09
0.95 5.41 (5.98) 1.55 0.56 (1.88) 0.53 4.85 (6.68) 1.67 0.27 (0.36) 0.07 0.85 (2.74) 0.59 -0.58 (-0.91) -0.18
0.99 5.18 (5.54) 1.52 0.71 (2.19) 0.61 4.47 (5.83) 1.62 0.17 (0.22) 0.04 0.95 (3.06) 0.65 -0.78 (-1.17) -0.24
Table C3: Moments value-weighted decile portfolios. This table shows average monthly returns for individual decile (1-10) and long-short value-weighted portfolios formed on the forecasted mean, median, volatility, skewness, and kurtosis. Results are reported for the full and liquid samples as well as individual regions, covering the period from 1995 to 2018. Values in brackets are t-statistics adjusted for Newey-West standard errors with 12 lags.
Mean
Full Sample Liquid Sample
Decile USA Europe Japan Asia Pacific USA Europe Japan Asia Pacific
1 -1.84 (-2.37) -2.18 (-3.20) -1.46 (-2.31) -3.27 (-4.24) -0.65 (-1.00) -0.40 (-0.88) -0.80 (-1.49) -0.88 (-1.37)
2 -0.74 (-1.24) -0.87 (-1.54) -0.64 (-1.15) -1.64 (-2.44) 0.28 (0.75) 0.37 (0.96) -0.24 (-0.47) 0.17 (0.34)
3 0.10 (0.22) -0.40 (-0.80) -0.37 (-0.77) -1.08 (-1.65) 0.60 (2.01) 0.44 (1.13) -0.09 (-0.21) 0.34 (0.67)
4 0.49 (1.42) 0.48 (1.17) -0.08 (-0.19) -0.59 (-0.95) 0.57 (1.96) 0.66 (1.92) 0.09 (0.22) 0.50 (1.21)
5 0.74 (2.22) 0.73 (1.91) 0.28 (0.68) 0.23 (0.46) 0.86 (3.26) 0.69 (2.02) 0.19 (0.48) 0.72 (1.74)
6 0.89 (2.84) 0.94 (2.74) 0.31 (0.76) 0.73 (1.41) 0.91 (3.45) 0.75 (2.17) 0.31 (0.90) 1.10 (2.78)
7 1.28 (3.91) 1.25 (3.43) 0.62 (1.52) 1.30 (2.45) 1.05 (3.71) 0.83 (2.43) 0.45 (1.26) 1.26 (3.87)
8 1.38 (4.07) 1.46 (3.84) 0.91 (2.31) 1.68 (3.34) 1.14 (3.83) 1.09 (3.22) 0.57 (1.55) 1.32 (3.25)
9 1.84 (4.51) 1.89 (4.35) 1.24 (2.59) 2.16 (3.91) 1.17 (3.51) 1.27 (3.17) 1.08 (3.09) 1.50 (3.73)
10 2.46 (5.05) 2.36 (4.86) 1.88 (3.46) 3.50 (5.42) 2.11 (5.32) 1.45 (3.51) 1.10 (2.50) 1.85 (4.01)
10-1 4.30 (6.63) 4.54 (7.38) 3.34 (8.10) 6.77 (15.25) 2.76 (5.02) 1.85 (5.26) 1.90 (6.07) 2.73 (6.18)
Median
Full Sample Liquid Sample
Decile USA Europe Japan Asia Pacific USA Europe Japan Asia Pacific
1 -1.84 (-2.11) -2.55 (-3.25) -1.46 (-1.89) -2.44 (-2.75) -0.53 (-0.75) -0.42 (-0.81) -0.87 (-1.45) -1.00 (-1.39)
2 -0.61 (-0.83) -1.07 (-1.75) -0.87 (-1.45) -1.42 (-2.04) 0.11 (0.25) 0.26 (0.61) -0.26 (-0.53) -0.01 (-0.02)
3 -0.21 (-0.34) -0.28 (-0.51) -0.27 (-0.49) -0.78 (-1.17) 0.48 (1.37) 0.48 (1.19) -0.12 (-0.25) 0.36 (0.68)
4 0.18 (0.34) 0.09 (0.24) -0.24 (-0.52) -0.83 (-1.27) 0.60 (2.11) 0.59 (1.69) -0.01 (-0.02) 0.44 (0.98)
5 0.50 (1.40) 0.44 (1.22) 0.03 (0.06) -0.34 (-0.55) 0.70 (2.46) 0.54 (1.56) 0.29 (0.70) 0.40 (1.00)
6 0.51 (1.52) 0.40 (1.07) -0.03 (-0.08) -0.08 (-0.16) 0.91 (3.73) 0.69 (2.16) 0.12 (0.36) 1.10 (2.75)
7 0.84 (3.00) 0.77 (2.31) 0.26 (0.64) 0.18 (0.38) 1.06 (4.01) 0.90 (2.62) 0.41 (1.09) 1.04 (3.00)
8 0.79 (2.58) 0.77 (2.51) 0.41 (1.03) 0.92 (1.94) 1.02 (3.39) 1.11 (3.24) 0.48 (1.46) 1.39 (3.71)
9 1.21 (4.17) 1.15 (3.01) 0.67 (1.66) 1.26 (2.94) 1.27 (4.59) 1.21 (3.27) 0.91 (2.52) 1.43 (3.76)
10 1.76 (5.06) 1.58 (3.90) 1.05 (2.49) 1.82 (3.87) 1.86 (4.94) 1.43 (3.58) 1.08 (2.69) 1.85 (4.31)
10-1 3.59 (5.02) 4.13 (6.86) 2.51 (4.86) 4.26 (7.67) 2.39 (4.43) 1.85 (4.42) 1.95 (5.10) 2.86 (5.70)
Volatility
Full Sample Liquid Sample
Decile USA Europe Japan Asia Pacific USA Europe Japan Asia Pacific
1 0.84 (4.03) 0.78 (3.24) 0.29 (1.41) 0.83 (3.03) 0.81 (4.01) 0.72 (3.17) 0.23 (0.97) 0.69 (2.50)
2 0.97 (3.61) 0.72 (2.45) 0.22 (0.73) 1.03 (2.46) 0.90 (3.75) 0.73 (2.47) 0.22 (0.68) 0.97 (2.50)
3 0.93 (3.03) 0.74 (2.25) 0.28 (0.77) 1.15 (2.60) 0.87 (3.18) 0.75 (2.37) 0.25 (0.57) 1.05 (2.47)
4 1.15 (3.44) 0.71 (1.85) 0.28 (0.75) 0.84 (1.55) 1.02 (3.20) 0.71 (1.94) 0.04 (0.09) 1.24 (2.51)
5 0.83 (1.88) 0.77 (1.77) 0.20 (0.53) 0.86 (1.56) 0.98 (3.00) 0.66 (1.61) 0.41 (1.18) 1.02 (2.12)
6 1.10 (2.24) 0.74 (1.50) 0.32 (0.65) 0.81 (1.28) 0.97 (2.38) 0.67 (1.66) 0.05 (0.13) 0.73 (1.44)
7 0.83 (1.56) 0.80 (1.46) 0.41 (0.76) 0.21 (0.30) 0.80 (1.67) 0.54 (1.19) 0.13 (0.25) 0.58 (1.04)
8 0.71 (1.09) 0.01 (0.01) 0.41 (0.69) -0.01 (-0.02) 1.00 (1.83) 0.58 (1.08) 0.46 (0.87) 0.79 (1.27)
9 0.64 (0.81) 0.05 (0.07) 0.66 (0.82) 0.06 (0.08) 0.69 (1.13) 0.81 (1.36) 0.53 (0.74) 0.78 (0.99)
10 0.16 (0.18) -0.87 (-0.93) -0.07 (-0.08) 1.07 (1.04) 0.32 (0.38) 0.13 (0.18) -0.22 (-0.25) -0.04 (-0.05)
10-1 -0.67 (-0.76) -1.65 (-1.93) -0.35 (-0.44) 0.24 (0.28) -0.49 (-0.63) -0.59 (-1.00) -0.45 (-0.59) -0.73 (-1.14)
Skewness
Full Sample Liquid Sample
Decile USA Europe Japan Asia Pacific USA Europe Japan Asia Pacific
1 0.79 (2.56) 0.65 (1.90) 0.22 (0.63) 0.78 (1.88) 0.55 (2.36) 0.33 (1.05) 0.06 (0.19) 0.78 (2.37)
2 1.01 (4.11) 1.05 (3.06) 0.35 (0.88) 1.13 (3.31) 0.72 (3.02) 0.44 (1.16) -0.23 (-0.59) 0.59 (1.48)
3 0.96 (2.95) 0.97 (2.81) 0.39 (0.82) 0.82 (1.74) 0.80 (2.88) 0.61 (1.87) 0.14 (0.33) 0.88 (1.98)
4 0.81 (1.85) 0.48 (1.37) 0.19 (0.38) 0.70 (1.32) 0.77 (2.58) 0.74 (2.22) 0.23 (0.53) 0.87 (1.92)
5 0.82 (1.70) 0.60 (1.37) 0.50 (0.96) 0.40 (0.72) 0.86 (2.62) 0.78 (2.09) 0.32 (0.76) 1.14 (2.86)
6 0.80 (1.56) 0.50 (0.99) 0.08 (0.15) 0.79 (1.31) 0.99 (2.63) 0.72 (2.29) 0.61 (1.44) 1.09 (2.41)
7 0.95 (1.68) 0.11 (0.20) 0.40 (0.74) 1.90 (1.22) 1.15 (2.98) 0.88 (2.23) 0.56 (1.31) 1.26 (2.46)
8 0.81 (1.66) 0.50 (1.07) 0.39 (0.74) 0.46 (0.70) 1.18 (3.03) 0.78 (1.95) 0.74 (1.71) 1.18 (2.62)
9 0.85 (1.61) -0.39 (-0.70) 0.42 (0.74) 0.32 (0.50) 1.31 (2.60) 1.10 (2.63) 0.60 (1.23) 1.20 (2.08)
10 0.89 (1.73) -0.26 (-0.50) -0.10 (-0.21) 0.35 (0.52) 0.75 (0.96) 0.71 (1.14) 0.31 (0.53) 0.83 (1.23)
10-1 0.10 (0.23) -0.91 (-2.32) -0.32 (-0.89) -0.44 (-0.95) 0.19 (0.27) 0.38 (0.78) 0.25 (0.60) 0.05 (0.09)
Kurtosis
Full Sample Liquid Sample
Decile USA Europe Japan Asia Pacific USA Europe Japan Asia Pacific
1 1.05 (2.54) 0.77 (2.08) 0.44 (0.95) 0.85 (1.92) 0.50 (0.90) 0.38 (0.61) 0.11 (0.15) 0.57 (0.92)
2 0.92 (3.29) 0.80 (2.31) 0.28 (0.84) 0.88 (2.18) 0.80 (1.97) 0.71 (1.81) 0.38 (0.85) 0.85 (1.13)
3 0.74 (2.70) 0.85 (2.86) 0.25 (0.75) 0.71 (1.99) 0.81 (2.99) 0.48 (1.24) 0.30 (0.66) 0.71 (1.78)
4 0.82 (3.10) 0.68 (2.38) 0.06 (0.16) 0.66 (1.58) 0.75 (2.67) 0.60 (1.68) 0.12 (0.32) 1.03 (2.56)
5 0.80 (2.54) 0.61 (1.52) 0.26 (0.64) 0.50 (1.17) 0.93 (3.30) 0.69 (1.92) 0.06 (0.15) 0.60 (1.34)
6 0.83 (2.48) 0.42 (1.11) 0.37 (0.74) 1.10 (1.93) 0.91 (3.12) 0.76 (2.32) 0.26 (0.78) 0.85 (2.13)
7 0.77 (2.05) 0.82 (1.89) 0.28 (0.67) 0.87 (1.69) 0.90 (3.01) 0.71 (2.02) 0.14 (0.39) 1.01 (2.49)
8 0.93 (2.70) 0.16 (0.56) 0.35 (0.80) 0.79 (1.57) 1.01 (3.02) 0.76 (2.27) 0.18 (0.46) 0.82 (1.95)
9 1.05 (3.06) -0.07 (-0.18) 0.23 (0.50) 0.75 (1.39) 1.11 (3.47) 0.81 (2.42) 0.40 (1.01) 1.21 (2.79)
10 0.94 (4.72) -0.24 (-0.62) -0.19 (-0.53) 0.51 (0.46) 0.80 (2.45) 0.55 (1.62) 0.22 (0.57) 1.02 (2.20)
10-1 -0.11 (-0.31) -1.01 (-2.72) -0.62 (-1.68) -0.34 (-0.29) 0.29 (0.76) 0.17 (0.44) 0.11 (0.24) 0.46 (1.40)
Appendix D Hyperparameter search and validation
We conduct a hyperparameter search using U.S. data from the period 1973 to 1994, employing weekly sampling and 22-day ahead returns, consistent with the setup for the main results in the paper. The period from 1973 to 1989 is utilized for training models with various hyperparameter values we are exploring. The period from 1990 to 1994 is used to evaluate the trained models and select the best hyperparameter values. Forecasts for the validation period are generated using the best hyperparameter values identified in the previous step. To maintain consistency with the train-test split used for the main results of the paper, we retrain the model annually during the validation period and use the most recent model for forecasting. Specifically, given a set of hyperparameter values, the model is initially trained on the 1973-1989 training sample and then rolled forward by one year four times, always using the model for the subsequent year’s forecasts. We utilize the full sample given the greater number of available observations compared to the liquid sample, which provides more stable results. Lengthening the validation sample would result in a shorter training sample, which is undesirable. The training sample should cover at least one severe market-wide negative shock in equities, which occurred in 1987. This approach suggests there is some potential room for improvement in finding better hyperparameters for the liquid sample and/or for international regions.
It is not feasible to perform a full hyperparameter search for all possible hyperparameters of the neural networks. Therefore, we divide the hyperparameter search into two sequential steps and focus only on a subset of hyperparameters. The rest of the hyperparameters are fixed and are based on the optimal hyperparameters found previously in the literature, e.g., Gu et al. (2020) or Tobek and Hronec (2021). Table D1 shows the fixed hyperparameters and their values.
Table D1: Fixed Hyperparameters. This table lists the hyperparameters that are fixed during the hyperparameter search. The number of epochs is set to $100*(A/n)$, where $n$ is the number of observations available for training and $A$ is an adjustment constant equal to $3{\small,}000{\small,}000$ for the full sample and $1{\small,}500{\small,}000$ for the liquid sample, derived from the average number of observations in individual samples over time.
Hyperparameter Value
Activation function LeakyReLU
Batch size 8192
Batch normalization True
Epochs $100*(A/n)$
Early-stopping patience 2 epochs
Early-stopping validation size 20%
Optimizer Adam
Adam momentum (0.9, 0.999)
Decay factor 1.0
Decay step size 100
L1 penalty for first layer 0.0001
Num. of networks in ensemble 20 liquid, 10 full sample
We start with a grid search for optimization algorithm hyperparameters, namely the learning rate (LR) and dropout rate (DR). We test values of $0.1$, $0.001$, and $0.0001$ for LR and 00, $0.2$, and $0.4$ for DR, evaluating the average quantile loss for the validation sample. We use a two-stage feedforward neural network with two hidden layers, each with $128$ neurons, as a starting point to perform the first round of hyperparameter search. The architecture was chosen to provide slightly higher capacity than neural networks used in Gu et al. (2020) or Tobek and Hronec (2021). Our evaluation criterion is the average quantile loss during the validation period from 1990 to 1994. Table D2 displays the results of the hyperparameter search.
Table D2: Hyperparameter Search for Learning Rate and Dropout. We conduct a hyperparameter search for the period from 1973 to 1994 using U.S. data only. We search for the optimal learning rate and dropout rate for the two-stage neural network. The rest of the hyperparameters are fixed as shown in Table D1. Loss is calculated as the average quantile loss for the validation period from 1990 to 1994.
Learning Rate Dropout Rate Loss
0.01 0.0 0.022827
0.01 0.2 0.022982
0.01 0.4 0.023135
0.001 0.0 0.022958
0.001 0.2 0.022821
0.001 0.4 0.022881
0.0001 0.0 0.022936
0.0001 0.2 0.022835
0.0001 0.4 0.022844
0.0003 0.2 0.022778
0.003 0.2 0.022914
0.0003 0.3 0.022815
0.0003 0.1 0.022829
The lowest average quantile loss is achieved for the learning rate of $0.0003$ and dropout rate of $0.2$, although dropout rates of $0.1$ and $0.3$ are not significantly worse than $0.2$. Fixing the learning rate and dropout rate at the optimal values achieved in the first part of the hyperparameter search, we perform the second part of the hyperparameter search to find an optimal architecture for the two-stage feedforward neural network, where we try different numbers of neurons and layers in the individual stages of the network. Table D3 shows the results of the architecture hyperparameter search.
Table D3: Hyperparameter Search for Architecture. We search for the optimal number of neurons and layers in the individual stages of the two-stage feed-forward neural network. The loss function is the average quantile loss achieved during the validation period from 1990 to 1994.
Stage1 (block) Stage2 Loss
2x128 8x1 0.022778
2x128 0x1 0.022815
2x128 16x1 0.022822
2x64 8x1 0.022829
2x256 8x1 0.022786
3x128 8x1 0.022810
1x128 8x1 0.022816
4x64 8x1 0.022872
A smaller or greater number of neurons than $8$ in the second stage leads to worse results. A smaller number of neurons is preferred to prevent potential overfitting and to speed up learning. The network already needs to be heavily regularized to achieve satisfactory performance; it, therefore, should not require more learning capacity by adding more neurons.
The L1 and L2 penalties were selected in isolation as it is not computationally possible to add them as another hyper-parameter in the search. The chosen values proved to be extremely robust and to work well across all the network architectures. The L1 penalty for the first layer of the first stage and L1/L2 penalty for the first layer of the second stage were selected based on inspection of estimated weights on the 1973-1994 training sample. The L1 penalty for the first layer of the first stage was selected using a one layer network and inspecting how many weights are non-negative. [^14] Note that one layer neural network is just a standard quantile linear regression with L1 penalty on its weights that is estimated using stochastic gradient descent. The goal was to induce some selection of input variables and to limit the effects of collinearity. The L1 penalty for the first stage is $0.0001$, and it forces weights of about $100-150$ of the input variables to zero. [^15] The L1 penalty for the first stage is changed to $0.00001$ for neural network specification, where raw returns are directly predicted from the first stage to reflect the smaller scale of the predicted variable.
The L1 and L2 penalties for the second stage were chosen solely to limit collinearity issues without any focus on variable selection. The penalty was chosen the smallest possible while still leading to no obvious collinearlity issues. [^16] The goal was to not have a mix of very large negative and positive weights, which are a symptom of the collinearity. The L1 and L2 penalties for the second stage are the same at $0.00001$. The L1 penalty for second stage is approximately $10$ times smaller than the L1 penalty for the first stage. The scale of the standardized stock returns, serving as a label for the first stage, is close to $1$. The scale of the raw stock returns, serving as a label for the second stage, is roughly $10$ times smaller. The L1 penalties for both stages are, therefore, consistent given the scale of the variables used as an input for the loss function.
D.0.1 Validation
Figure D.1 illustrates the data split into training, validation, and testing sets for the U.S. and international samples. Hyperparameter optimization and model training are conducted solely on U.S. data, using international data only for out-of-sample evaluation.
[…]
Figure D.1: Validation sample-splitting schema. This figure shows how the data is split into training ($m_{i}$), validation ($hpo$) and testing ($p_{i}$) sets for the U.S. and international samples. All models are trained on the U.S. data and then used to generate out-of-sample predictions for both the U.S. and the international sample. Models are retrained each year using an expanding window of data reaching back to 1973. Predictions are generated monthly, always using the most up-to-date model available at the end of each month, e.g. $p_{1}$ predictions represent all monthly predictions during the period of year $1$ and are generated using the model trained on $m_{1}$.
In order to maximize the number of observations available for training, future returns are computed each week along with all the required inputs to the neural network. Future returns cover the next 22 business days [^17] The count of business days here also includes holidays for simplicity., approximately one month of trading. This leads to approximately 4.3 times more partially overlapping observations available for training compared to using non-overlapping monthly data only [^18] Using only monthly data during the training leads to inferior models because it can miss some reverting rare price changes.. Weekly sampled future returns are used only during the model training. Out-of-sample predictions are generated and evaluated using regular calendar months.
Appendix E Algorithms for approximating probability distribution function and moments from quantiles
Algorithm 1 contains the algorithm for deriving the probability density function from the quantiles. The algorithm is referred to as quantiles-to-density in the paper. Algorithm 2 contains the algorithm for deriving the moments of the probability distribution function. The algorithm is referred to as density-to-moments in the paper, and the algorithm below does not contain the adjustment of the moments described in Section E.3.
E.1 Fallback to linear B-Splines
The liquid universe of stocks generally exhibits well-behaved density functions. Therefore, the fallback to the linear approximation of the density is rarely needed. However, it is sometimes necessary for the full sample due to the presence of infrequently traded stocks with small capitalization. There can be many days with zero trading activity for micro-cap stocks, especially in the historical period. Days with zero trading are a well-documented proxy for liquidity in the finance literature, formally introduced by Lesmond et al. (1999). See Goyenko et al. (2009) for an overview of other related liquidity proxies along with a comparison of their relative performance. This method could also be applied in bond returns forecasting where high illiquidity and zero-return observations are common. The presence of zero-return days can lead to a large probability mass around zero and, in extreme cases, to a discrete probability of a zero return. The continuous density function implied from the cumulative distribution function is then not well-behaved, and the fallback to linear approximation becomes necessary.
For larger cap stocks where the density fallback was applied, the cause was manually checked. The density always exhibited a large mass around zero, indicating that the model was predicting a discrete probability of zero return in the following month. Fallback predictions for large cap stocks were often associated with a corporate event, such as being close to official delisting or a recent trading halt. One such case is detailed in Figure E.1, depicting forecasts for October 2014 for Idenix Pharmaceuticals, which was delisted just five days later. Idenix Pharmaceuticals received a tender takeover offer from Merck on June 9, 2014191919The announcement of the acquisition is available here https://www.merck.com/news/merck-to-acquire-idenix/..
Figure E.1: Example of density with a large mass around zero. This figure shows the derived probability density function for monthly return predictions for Idenix Pharmaceuticals in October 2014.
Algorithm 1 PDF
[…]
2:$gp=100$triangleright Density of grid points
3:$d_{min}=1e-5$triangleright Minimum value for density function
$eps=1e-4$triangleright Epsilon
5:$z=0$triangleright Keep only the highest quantile with -1 return
6:for $i$ in 1 to length(x) - 1 do
7:    if $x[i+1]=-1$ then
8:        $z=z+1$
9:    end if
10:end for
[…]
13:if $min(y)<0.05$ then $\triangleright$ Get 0.9 and 0.1 quantiles for outlier detection
14:    $x_{Q10}=x[y=0.1]$
15:else
16:    $x_{Q10}=x[3]$
17:end if
18:$x_{Q90}=x[y=0.9]$
19:for $i$ in 1 to length(x) - 1 do $\triangleright$ Fix inconsistent quantile predictions
20:    if $x[i+1]<x[i]+eps$ then
21:        $x[i+1]=x[i]+eps$
22:    end if
23:end for
24:allocate empty vector $\tilde{\textbf{x}}$triangleright Create grid of denser x values
25:for $i$ in 1 to length(x) - 3 do
26:    $x_{l}=x[i+1]$
27:    $x_{u}=x[i+2]$
28:    add $range(x_{l},x_{u},(x_{u}-x_{l})/gp)$ to $\tilde{\textbf{x}}$
29:end for
30:drop duplicates in $\tilde{\textbf{x}}$
31:sort values of $\tilde{\textbf{x}}$ from the smallest to the largest
32:fit cubic B-spline interpolation $S$ on x and y
33:use $S$ to fit its first derivative on $\tilde{\textbf{x}}$: $\textbf{d}=S^{\prime}(\tilde{\textbf{x}})$
34:if $min(\textbf{d}[(\tilde{\textbf{x}}\geq x_{Q10})\&(\tilde{\textbf{x}}\leq x_{Q90})])<d_{min}$ then $\triangleright$ Density fallback
35:    fit linear B-spline interpolation $S$ on x and y
36:    use $S$ to fit its first derivative on $\tilde{\textbf{x}}$: $\textbf{d}=S^{\prime}(\tilde{\textbf{x}})$
37:end if
38:d[$\textbf{d}[\textbf{d}<d_{min}]=d_{min}$triangleright Enforce minimum density
39:$w=min(\textbf{d}[(\tilde{\textbf{x}}<=x_{Q10})])$triangleright Enforce monotonicity in the tails
40:$v=argmin(\textbf{d}[(\tilde{\textbf{x}}<=x_{Q10})])$
[…]
42:$w=min(\textbf{d}[(\tilde{\textbf{x}}>=x_{Q90})])$
43:$v=argmin(\textbf{d}[(\tilde{\textbf{x}}>=x_{Q90})])$
[…]
45:$x_{min}=min(x)$triangleright get max and min of x
46:$x_{max}=max(x)$
47:if $min(\tilde{\textbf{x}})<-1$ then $\triangleright$ Truncate the distribution at -1 if needed
48:    $\textbf{d}=\textbf{d}[\tilde{\textbf{x}}\geq-1]$
49:    $\tilde{\textbf{x}}=\tilde{\textbf{x}}[\tilde{\textbf{x}}\geq-1]$
50:    $x_{min}=-1$
51:end if
52:$P_{l}=1-S(x_{max})$triangleright Compute discrete probabilities at tails
53:$P_{u}=S(x_{min})$
E.2 Central Moments
We can derive central moments from the non-central moments using the Equation 9.
$\sigma$ $=m_{2}-m_{1}^{2}$ (9)
skewness $=\frac{m_{3}-3m_{1}\sigma-m_{1}^{3}}{\sigma^{3/2}}$
kurtosis $=\frac{m_{4}-4m_{1}m_{3}+6m_{1}^{2}m_{2}-3m_{1}^{4}}{\sigma^{2}}$
where $m_{i}$ for $i\in{1,2,3,4}$ are non-central moments and $\sigma$ is the variance.
E.3 Adjustment of moments
The required moment adjustments are estimated using simulated data with known precise theoretical moments. One million non-central Student’s t distributions are generated with various degrees of freedom and noncentrality parameter.  [^20] Non-central Student’s t distribution is defined as $\frac{Y+nc}{\sqrt{V/df}}$ where $Y$ is a standard normal distribution, $V$ is an independent chi-square random variable, $nc$ is noncentrality parameter, and $df$ are degrees of freedom. See scipy.stats.nct documentation for more details on its implementation. Degrees of freedom are randomly selected from {5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 30, 40, 50, 60, 70, 80, 90, 100, 1000, 10000} with replacement and equal probability of each outcome. The noncentrality parameter is randomly selected using uniform distribution of $[-0.5,5]$ interval focusing mostly on positive skewness, as is observed in the empirical equities data. The scale of the generated distributions is set to 0.1 to limit the effects of the fact that stock returns cannot be smaller than $-1$, which is reflected in the density-to-moments algorithm but not in the theoretical distributions. Precise theoretical moments of the distribution are saved, and quantiles of the distribution are fed into the “naive“ density-to-moments algorithm to compute predicted moments. Generated distributions with theoretical kurtosis larger than $20$ are discarded as they would create strong leverage points and skew the estimated parameters to capture these extremes.
The values of the parameters are then estimated using OLS, where precise theoretical moments are predicted using moments produced by the “naive“ density-to-moments algorithm. The simple linear regression is preferred here as it provides a satisfactory fit while being easy to interpret and more likely to generalize to other distributions. [^21] It is possible to further improve the fit by employing neural networks or higher order polynomials and interactions between the exogenous variables if required.
The following adjustment is proposed to decrease bias in the moments estimated from the “naive“ density-to-moments algorithm.
$k_{e}$ $=k-3$ (10)
$\tilde{v}$ $=v\times(1.0023-0.0021\times s+0.0022\times k_{e})$
$\tilde{s}$ $=0.9950\times s+0.0261\times s^{2}+0.0107\times k_{e}$
$\tilde{k}$ $=3+1.4185\times k_{e}+0.0466\times k_{e}^{2}-0.7395\times s$
Where $v,s,k$ are variance, skewness, and kurtosis from the “naive“ density-to-moments algorithm, $\tilde{v},\tilde{s},\tilde{k}$ are their adjusted counterparts, and $k_{e}$ is excess kurtosis. All of the estimated parameters are significant with t-distribution over $100$ in absolute terms. The fit is high for all the regressions with $R^{2}$ over $0.9$.
Variance $v$ requires only a small adjustment, which mainly accounts for the fact that it is underestimated for heavy-tailed distributions with high kurtosis. The adjustment for skewness $s$ is larger to account for its underestimation when either skewness or kurtosis of the distribution is high. Finally, as expected, kurtosis requires the largest adjustment due to the difficulty of capturing it with the “naive“ density-to-moments algorithm.
The adjusted density-to-moments algorithm is next tested on probability distributions that are close to those actually observed in the predictions of the future return distribution. The predicted densities from neural networks have positive skewness and are heavy-tailed with positive excess kurtosis. A range of distributions is tested in Table E1. Normal distribution is the starting point due to its wide use in quantitative finance. t-distribution with small degrees of freedom then serves as a benchmark for a more heavy-tailed distribution. Finally, non-central Student’s t-distribution is the closest match to the distribution actually observed in the data.
Table E1: Adjustment fit of density-to-moments algorithm. $m,v,s,k$ denote mean, variance, skewness, and kurtosis estimated using the “naive“ (non-adjusted) density-to-moments algorithm defined in Section E.3, $m_{t},v_{t},s_{t},k_{t}$ corresponding theoretical values and $\tilde{v},\tilde{s},\tilde{k}$ are generaged using adjusted density-to-moments algorithm, i.e. values are adjusted using equations 10.
Dist $m_{t}$ $m$ $v_{t}$ $v$ $\tilde{v}$ $s_{t}$ $s$ $\tilde{s}$ $k_{t}$ $k$ $\tilde{k}$
Normal $0$0.000 1.000 0.998 1.001 $0$$0$0.000 3.000 2.980 2.971
t: df=10 $0$0.000 1.250 1.244 1.250 $0$0.000 0.009 4.000 3.852 4.242
t: df=6 $0$0.000 1.500 1.487 1.497 $0$0.000 0.023 6.000 5.161 6.282
t: df=5 $0$0.000 1.667 1.646 1.662 $0$0.001 0.035 9.000 6.360 8.293
nct: df=5, nc=1 0.119 0.119 1.919 1.892 1.914 1.266 1.116 1.200 13.321 8.385 11.164
nct: df=6, nc=3 0.345 0.345 3.072 3.035 3.076 1.832 1.686 1.823 12.991 9.691 13.330
nct: df=5, nc=4 0.476 0.476 5.698 5.593 5.736 2.718 2.372 2.644 29.832 15.827 27.104
Table E1 documents a good fit for lower moments and even for kurtosis post-adjustment. The computed mean is almost identical to the theoretical mean, as is the computed variance. Skewness benefits from the moment adjustment for heavy-tailed positively skewed distribution while exhibiting reasonable fit even without the adjustment. Kurtosis is the hardest to fit precisely, but the adjustment is quite precise even for the cases with the highest theoretical kurtosis. This section has, therefore, confirmed that the proposed moment estimation technique from the quantiles of the distribution yields reasonable results for empirical analysis.
Algorithm 2 Moments
1:$\textbf{x},\textbf{y},P_{l},P_{u},x_{min},x_{max}$
2:$m_{0},m_{1},m_{2},m_{3},m_{4}=0,0,0,0,0$
3:for $i$ in 1 to length(x) - 1 do
4:    $x_{1}=\textbf{x}[i]$
5:    $x_{2}=\textbf{x}[i+1]$
6:    $y_{1}=\textbf{y}[i]$
7:    $y_{2}=\textbf{y}[i+1]$
8:    $b=(y_{2}-y_{1})/(x_{2}-x_{1})$
9:    $a=y_{1}-bx_{1}$
10:    $m_{0}=m_{0}+ax_{2}+1/2bx_{2}^{2}-ax_{1}-1/2bx_{1}^{2}$
11:    $m_{1}=m_{1}+1/2ax_{2}^{2}+1/3bx_{2}^{3}-1/2ax_{1}^{2}-1/3bx_{1}^{3}$
12:    $m_{2}=m_{2}+1/3ax_{2}^{3}+1/4bx_{2}^{4}-1/3ax_{1}^{3}-1/4bx_{1}^{4}$
13:    $m_{3}=m_{3}+1/4ax_{2}^{4}+1/5bx_{2}^{5}-1/4ax_{1}^{4}-1/5bx_{1}^{5}$
14:    $m_{4}=m_{4}+1/5ax_{2}^{5}+1/6bx_{2}^{6}-1/5ax_{1}^{5}-1/6bx_{1}^{6}$
15:end for
16:$m_{1}=m_{1}+P_{l}x_{min}+P_{u}x_{max}$triangleright Add discrete probabilities in the tails
17:$m_{2}=m_{2}+P_{l}x_{min}^{2}+P_{u}x_{max}^{2}$
18:$m_{3}=m_{3}+P_{l}x_{min}^{3}+P_{u}x_{max}^{3}$
19:$m_{4}=m_{4}+P_{l}x_{min}^{4}+P_{u}x_{max}^{4}$
20:$m_{1}=m_{1}/m_{0}$triangleright Normalize probability measure to 1
21:$m_{2}=m_{2}/m_{0}$
22:$m_{3}=m_{3}/m_{0}$
23:$m_{4}=m_{4}/m_{0}$
24:$variance=m_{2}-m_{1}^{2}$
25:$skewness=(m_{3}-3m_{1}var-m_{1}^{3})/var^{3/2}$
26:$kurtosis=(m_{4}-4m_{1}m_{3}+6m_{1}^{2}m_{2}-3m_{1}^{4})/var^{2}$